### v2

# Semiclassical measure of the propagation between two topological insulators

Abstract
We study the propagation of a family of initial conditions in the presence of two topological insulators without magnetic field where the interface is a smooth connected curve not compact without boundaries. The solution is governed by an adiabatic modulation of a Dirac operator with a variable mass. We determine the evolution of the semiclassical measure of the solution with a two-scale Wigner measure method by reducing the Dirac operator to a normal form.
Keywords : propagation, Dirac operator, semiclassical analysis, Wigner measure, semiclassical measure valued in a separable Hilbert space, time dependent two-scale semiclassical measure.
$2024$ Mathematics Subject Classification : 35B40, 35F05, 35Q40.
Acknowledgments. The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions. The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.
1 Introduction
1.1 Topological insulators and Dirac type equations
Topological insulators are materials that are insulating in their bulk but support robust currents along their boundaries, meaning that electrons can only move along the surface of the material. The Dirac equation has been derived in the context of topological insulators for honeycomb structures [9, 8, 5] such as graphene [22, 18, 7]. From a global point of view, Dirac-type equations are often a simpler continuum macroscopic description of transport in a narrow energy band near the band crossing [2, 12, 24].
In this paper we consider the following system of the Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
$\left\{\begin{array}[]{ll}\Big{(}\varepsilon\mathrm{D}_{t}+\mathrm{m}(x)\sigma_{3}+\varepsilon\mathrm{D}_{1}\sigma_{1}+\varepsilon\mathrm{D}_{2}\sigma_{2}\Big{)}\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right.$ (1)
where ${\varepsilon}$ is a small positive parameter, called semiclassical parameter. The function $\psi^{\varepsilon}(t,\cdot)$ also denoted $\psi^{\varepsilon}_{t}$ is interpreted as the wave function of an electron at time $t$ in a material consisting in two topological insulators. The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is a uniformly bounded (in $\varepsilon$) family of ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ of initial wave functions of the electron. The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function (i.e. $\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and all its derivatives are bounded. The matrices ${\left(\sigma_{j}\right)_{1\leqslant j\leqslant 3}}$ are the Pauli matrices
$\sigma_{1}\coloneqq\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\;,\;\sigma_{2}\coloneqq\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}\;,\;\sigma_{3}\coloneqq\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
We shall use the notation
$\mathrm{H}_{\varepsilon}\coloneqq\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}.$
Then the equation (1) rewrites $\varepsilon\mathrm{D}_{t}\psi_{t}^{\varepsilon}+\mathrm{H}_{\varepsilon}\psi_{t}^{\varepsilon}=0$.
In this model, the function $\mathrm{m}$ splits the space into two parts, which are the two topological insulators. From a modeling point of view, the function $\mathrm{m}$ is positive on one of the topological insulators and negative on the other so the interface between the two topological insulators is defined by
$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\ |\ \mathrm{m}(x)=0\right\}.$
We assume the set $\mathrm{E}$ is a connected curve not compact without boundaries. From a physical point of view, the sign of $\mathrm{m}$ is the insulating phase of a material. This generates a unidirectional propagation along the interface $\mathrm{E}$. This propagation is one of the most important issue of physical applications of topological insulators and superconductors [2, 24], photonics [19, 22, 23], acoustics [21, 20] and fluid mechanics [13].
Our aim is to describe the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to equation (1) for initial conditions in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. At least, we can understand limits as $\varepsilon$ goes to 00 by constructing semiclassical defect measures, also called Wigner measures. Then, we can deduce various properties of this kind of measures and thereby indirectly recover information about asymptotics as $\varepsilon$ goes to 00.
1.2 Edge states generated by coherent states
The system (1) has already been studied in [1] where are performed precise computation. The edge mode computed is generated by a very specific shape of semiclassical wave packet.
In [6], it was proved that some terms of the solution to equation (1) are negligible in $\mathrm{L}^{\infty}$-norm as $\varepsilon$ goes to 00. It has been proven, for initial conditions, that semiclassical wave packets concentrate at a point on the $\mathrm{E}$ curve with no constraint on orientation. However, these dispersive terms are not necessarily negligible in the $\mathrm{L}^{2}$ norm. Let us recall these results.
Let $x_{0}\in\mathrm{E}$ and $t\mapsto x_{t}$ the solution to the ODE
$\left\{\begin{array}[]{l}\dot{x}_{t}=\frac{\nabla\mathrm{m}(x_{t})^{\perp}}{\left|\nabla\mathrm{m}(x_{t})\right|},\\ \left.x_{t}\right|_{t=0}=x_{0}.\end{array}\right.$
Let us define, for all $t\in\mathbb{R}$, $r_{t}\in\mathbb{R}_{>0}$ and $\theta_{t}\in\mathbb{R}$, such that $r_{t}\begin{pmatrix}-\sin\theta_{t}\\ \cos\theta_{t}\end{pmatrix}=\nabla\mathrm{m}(x_{t})$ and
$\vec{V}_{\theta_{t}}\coloneqq\begin{pmatrix}e^{\frac{-\mathrm{i}\theta_{t}}{2}}\\ -e^{\frac{\mathrm{i}\theta_{t}}{2}}\end{pmatrix}.$
Following [6], we define wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$ as a $\varepsilon$-dependent function satisfying for all positive $\varepsilon$ and for all $x\in\mathbb{R}^{2}$,
$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
with $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$. The parameter $\xi_{0}$ is interpreted as the impulsion (or the speed) of an electron of wave function $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$.
Theorem 1.1 ([1] Theorem 2 and [6] Theorem A.1).
Let $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then there exists $T>0$, such that for all $t\in(0,T)$, uniformly for $\varepsilon\in(0,1]$,
$\psi_{t}^{\varepsilon}=\mathrm{WP}^{\varepsilon}_{x_{t},0}+\mathcal{O}_{\mathrm{L}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{-1/4}\right)+\mathcal{O}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{1/2}\right).$ (2)
where $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ is a wave packet concentrated at $(x_{t},(0,0))$, oriented along $\vec{V}_{\theta_{t}}$. Moreover, we have the following.
  1. (i)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}$, then the equality (2) holds for all $t>0$ without $\mathcal{O}_{\mathrm{L}^{\infty}}\left(\varepsilon^{-1/4}\right)$.
  2. (ii)
    If there exists $f\in\mathscr{S}\left(\mathbb{R},\mathbb{C}\right)$ such that for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
    $\psi_{0}^{\varepsilon}(x)\coloneqq\frac{r_{0}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{0}\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{0}},$
    where $\mathrm{R}_{\theta}$ is the rotation of angle $\theta$ in $\mathbb{R}^{2}$, then uniformly for $\varepsilon\in(0,1]$, for all $(t,x)\in\mathbb{R}\times\mathbb{R}^{2}$,
    $\psi_{t}^{\varepsilon}(x)=\frac{r_{t}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{t}\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{t}}\ +\ \mathcal{O}\left(\varepsilon^{1/2}\langle t\rangle\right).$
  3. (iii)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}^{\perp}$, then $\mathrm{WP}^{\varepsilon}_{x_{t},0}=0$.
The initial wave packet is concentrated on the curve $\mathcal{C}\coloneqq\mathrm{E}\times\{(0,0)\}\subset\mathbb{R}^{4}$ and so is the wave packet at time $t$. Physically, the wave packets under consideration have no impulse.
Let us informally describe these wave packets. In the point (ii), the variables $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}$ and $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}$ represent the tangent and the normal variable to the curve $\mathrm{E}$, respectively. They do not play the same role in the considered wave packet. In particular, for all $t\in[0,T)$, the wave packet $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ have a Gaussian behavior in the normal variable and is localized near the point $(x_{t},(0,0))$ in the phase space.
If the orientation is not along $\vec{V}_{\theta_{0}}$ or the decay with respect to the normal variable is different, then the $\mathrm{L}^{\infty}$-remaining part is not necessarily negligible in $\mathrm{L}^{2}$-norm. Therefore Theorem 1.1 does not give a full description of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
To prove this theorem, the Hamiltonian $\mathrm{H}_{\varepsilon}$ is reduced to a normal form through Fourier Integral Operator in [6]. In the same way, we will reduce it through a change of variable to have a model problem which can be understand as if the curve $\mathrm{E}$ was a line. Our objective is two-fold: we aim at describing the evolution of these dispersive terms and we want to consider initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
1.3 Main result
Instead of computing the asymptotics of the wave function itself, we focus on the probability densities such as the position one defined by $\left|\psi^{\varepsilon}_{t}(x)\right|_{\mathbb{C}^{2}}^{2}\mathrm{d}x$. Its physical interpretation is the probability of finding the particle at time $t$ in the position $x$. This approach generalizes the previous strategy and catches the information on the quantities that are not negligible in $\mathrm{L}^{2}$-norm. More precisely, we want to describe the evolution as $\varepsilon$ goes to 00 of the quadratic quantities
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\coloneqq\int_{\mathbb{R}^{2}}\left\langle a(x)\psi^{\varepsilon}_{t}(x),\psi^{\varepsilon}_{t}(x)\right\rangle_{\mathbb{C}^{2}}\mathrm{d}x,$ (3)
where $\mathbb{C}^{2,2}$ is the set of $2$ by $2$ matrices and the scalar product of $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$ is $\left\langle v,w\right\rangle_{\mathbb{C}^{2}}\coloneqq v\cdot\overline{w}$.
For example, if $a$ is equal to a projector on a direction $\mathbb{R}\vec{V}$ for $\vec{V}\in\mathbb{R}^{2}\setminus\{0\}$ on a subspace $\Omega$, then $\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ describes the probability that $\psi^{\varepsilon}_{t}$ is collinear to $\vec{V}$ at time $t$ on the subspace $\Omega$.
We now introduce our assumptions on the function $\mathrm{m}$. Physically, the assumptions describe the interface between the two topological insulators.
Assumption 1.2.
The function $\mathrm{m}$ satisfies
$\inf_{x\in\mathrm{E}}{\Big{|}\nabla\mathrm{m}(x)\Big{|}}>0.$
To simplify the presentation and technical details, we will work with a global tubular neighborhood of the curve $\mathrm{E}$. To ensure the existence of such a neighborhood, we assume the following assumption. Actually, if we consider a neighborhood of a point of the curve $\mathrm{E}$, the local version of the following assumption will always be satisfied.
Let $x_{0}\in\mathrm{E}$. Let us consider the parametrization $\gamma$ of the curve $\mathrm{E}$ and the vector-valued function $\nu$ such that $\gamma(0)=x_{0}$ and for all $s\in\mathbb{R}$,
$\gamma^{\prime}(s)\coloneqq\frac{-{\nabla\mathrm{m}(\gamma(s))}^{\perp}}{\left|\nabla\mathrm{m}(\gamma(s))\right|}\;\;,\;\;\nu(s)\coloneqq\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|},$
where denotes the $\pi/2$-counterclockwise rotation. Then the curvature of the curve $\mathrm{E}$ at a point $\gamma(s)$ is the real number $\kappa(s)$ given by Frenet’s formula
$\kappa(s)\coloneqq\gamma^{\prime\prime}(s)\cdot\nu(s).$
Let us denote ${\mathrm{I}\coloneqq\left(-\frac{1}{2||\kappa||_{\infty}},\frac{1}{2||\kappa||_{\infty}}\right)}$.
Assumption 1.3.
The map
$\begin{matrix}\Phi:&\mathbb{R}\times\mathrm{I}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$ (4)
is a global diffeomorphism from ×$\mathbb{R}\times\mathrm{I}$ to
$\Omega\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}\right\}.$
[…]
$\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|}=\begin{pmatrix}\cos\theta(s)\\ \sin\theta(s)\end{pmatrix}.$
We will denote the dual variable of $(s,y)$ by $(\sigma,\eta)$. Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;\;,\;\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n||\nabla\mathrm{m}(\gamma(s))|}.$ (5)
Although the function $\lambda_{0}$ does not depend on $s$, we also will denote it by $\lambda_{0}(s,\sigma)$.
A consequence of our main result (in Section 3.3) is the following asymptotic description of the position density of the solution of equation (1) for initial data which are wave packets concentrating on the curve $\mathcal{C}$.
Theorem 1.4.
Assume Assumption 1.2 and 1.3. Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then, there exist a vanishing sequence ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ of positive numbers and a family of measurable maps ${t\mapsto\left(\mathrm{d}\gamma^{t}_{n}\right)_{n\in\mathbb{Z}}}$ valued in the set of scalar non negative Radon measures on $\mathbb{R}^{2}$, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\Xi(t)\left\langle a\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s)\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t$
with for all $s\in\mathbb{R}$,
$\mathrm{d}\rho_{t}(s)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
where for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\mathrm{d}\gamma_{n}^{t}&=&\partial_{\sigma}\mathrm{d}\gamma_{n}^{t}\ \partial_{s}\lambda_{n}-\partial_{s}\mathrm{d}\gamma_{n}^{t}\ \partial_{\sigma}\lambda_{n}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$ (6)
and $\mathrm{d}\gamma_{n,0}$ is determined by the initial condition.
We will see in Section 3.3 a more general result (Theorem 3.10), with initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
Remark (Link between Theorem 1.1 and Theorem 1.4).
With the notation of Theorem 1.4, the measure $\mathrm{d}\gamma_{0}^{t}$ is the measure associated to the part $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ of the solution of Theorem 1.1. Likewise, the measures $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ are associated to the $\mathrm{L}^{\infty}$-remainder part of the solution of Theorem 1.1. Notably, if for all $n\in\mathbb{Z}_{\neq 0},\mathrm{d}\gamma_{n,0}=0$ and $\mathrm{d}\gamma_{0,0}=1$, then for all $(t,s,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\mathrm{d}\gamma_{0}^{t}(s,\sigma)=\delta(s-t)\delta(\sigma)\quad,\quad\mathrm{d}\gamma_{n}^{t}(s,\sigma)=0,$
where $\delta$ is the Dirac mass over 00.
Theorem 1.4 gives a description of the evolution of the position density when initial data are wave packets. But Theorem 3.10 allow us to consider any initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ concentrating near the curve $\mathcal{C}$ to describe their asymptotic position density behavior. A consequence of this theorem is the following one.
Theorem 1.5.
Assume Assumption 1.2 and 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition, then,
$\mathrm{d}\nu\mathds{1}_{\mathrm{E}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathrm{E}}\mathrm{d}t$
where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6).
This theorem gives a lower bound estimation of the position density at time $t$ of the evolution of any initial data concentrating over the curve $\mathrm{E}$.
1.4 Main ideas of the proofs and organisation of the article
There are two main ideas in the proof of Theorem 3.10.
The first step is to transform the operator $\mathrm{H}_{\varepsilon}$ into a model operator as if the curve $\mathrm{E}$ were exactly a line. Let us discuss the case of the line.
If the interface is a line defined by a mass $\mathrm{m}$ such that for all $x\coloneqq(x_{1},x_{2})\in\mathbb{R}^{2},\mathrm{m}(x)=\widetilde{m}(x_{1})x_{2}$, then the operator $\mathrm{H}_{\varepsilon}$ is unitary equivalent to the following one,
$\sqrt{\varepsilon}\begin{pmatrix}\sqrt{\varepsilon}\mathrm{D}_{1}&x_{2}\widetilde{m}(x_{1})+i\mathrm{D}_{2}\\ x_{2}\widetilde{m}(x_{1})-i\mathrm{D}_{2}&-\sqrt{\varepsilon}\mathrm{D}_{1}\end{pmatrix}.$
This operator is $\sqrt{\varepsilon}$-semiclassical with respect to $x_{1}$ only. Thus, it will be considered as the quantization of an operator-valued symbol. It leads us to the second idea of the proof.
The second step is to build a two-scale quantization that fits the scale of our new operator. In fact, our analysis relies on a microlocal approach, replacing densities $\left(\left|\psi_{t}^{\varepsilon}(x)\right|^{2}\mathrm{d}x\right)_{\varepsilon>0}$ by the Wigner transform [26, 15]. The Wigner transform provides us with a description of the dynamics for a very general class of equations and initial data, allowing us to deal with quantum observables. Therefore, we will study the quadratic quantity (3), with an observable $a$ depending of position and momentum, i.e.
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\left(\cdot{},\varepsilon\mathrm{D}_{x}\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
We recall basic facts about these notions in Section 2. However, to precisely analyze what happens near the curve $\mathcal{C}$, we have to take into account how $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ concentrates on the curve $\mathcal{C}$ with respect to the scale $\sqrt{\varepsilon}$. We introduce in Section 3 the two-scale Wigner measure that be of crucial use for proving our main results Theorems 3.8 and 3.10. Finally, we prove Theorems 1.41.53.8 and 3.10 in Section 4.
1.5 Notation
We denote by $\mathbb{S}^{2}$ the unit sphere of $\mathbb{R}^{3}$.
We denote by $v\otimes w$ the $2$ by $2$ matrix $v\,^{t}w$ for $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$.
We denote by ${\rm tr}$ the trace in $\mathcal{L}\left(\mathbb{C}^{2}\right)$ and by $\mathrm{Tr}$ the trace in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)\right)$.
Let $\mathcal{H}$ a separable Hilbert space and $n$ an integer, for $f,g\in\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote the usual inner product by
$\left\langle f,g\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)}\coloneqq\int_{\mathbb{R}^{n}}\left\langle f(x),g(x)\right\rangle_{\mathcal{H}}\,\mathrm{d}x.$
For $A$ an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote by $A^{\dagger}$ the formal adjoint operator of $A$ associated to the usual scalar product of $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$.
We define the semiclassical Fourier transform for all $\varepsilon>0$ and for $f\in\mathrm{L}^{1}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ by
$\mathcal{F}_{\varepsilon}f(\xi)\coloneqq\frac{1}{2\pi\varepsilon}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\frac{x\cdot\xi}{\varepsilon}}f(x)\mathrm{d}x,\quad\xi\in\mathbb{R}^{2},$
and we denote by $\widehat{f}$ the Fourier transform of $f$ defined by $\widehat{f}\coloneqq\mathcal{F}_{1}f$.
2 Semiclassical analysis
Our problem requires an analysis in the phase space. We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1) and its weak limits in the space of distributions that are called Wigner measures. We recall the definition and the main properties of the Wigner transform in Section 2.1 below. Then, we apply these results to the solution of our problem in Section 2.2.
2.1 Wigner transform and Wigner measure
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. The Wigner transform of ${(f^{\varepsilon})_{\varepsilon>0}}$ is defined as follow, for all $(x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$,
$\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\coloneqq\int_{\mathbb{R}^{2}}e^{i\xi\cdot y}f^{\varepsilon}\left(x+\frac{\varepsilon y}{2}\right)\otimes\overline{f^{\varepsilon}\left(x-\frac{\varepsilon y}{2}\right)}\frac{\mathrm{d}y}{(2\pi)^{2}}.$
It was introduced by Wigner in [25] and in [26]. One can check that $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is valued in the set of Hermitian matrices and $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]\in\mathrm{L}^{2}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\cap\mathscr{C}^{0}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$. Moreover, one recovers formally the density in position through the relation
$\int_{\mathbb{R}^{2}}{\rm tr}\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\mathrm{d}\xi=\left|f^{\varepsilon}(x)\right|_{\mathbb{C}^{2}}^{2}.$
The function $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is usually considered as a distribution on the phase space $\mathbb{R}^{4}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$. Moreover, we have for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{4},\mathbb{C}^{2,2})$,
$\mathrm{I}^{\varepsilon}_{f}(a)\coloneqq\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi=\left\langle\mathrm{Op}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
where $\mathrm{Op}_{\varepsilon}(a)$ is the semiclassical pseudo-differential operator of symbol $a$ obtained by the Weyl quantization
$\mathrm{Op}_{\varepsilon}(a)f(x)\coloneqq\int_{\mathbb{R}^{4}}a\left(\frac{x+y}{2},\xi\right)f(y)e^{\mathrm{i}\frac{\xi\cdot(x-y)}{\varepsilon}}\frac{\mathrm{d}y\mathrm{d}\xi}{(2\pi\varepsilon)^{2}},\quad f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),\ x\in\mathbb{R}^{2}.$
The Calderón-Vaillancourt Theorem [4] asserts that the family of operators $\left(\mathrm{Op}_{\varepsilon}(a)\right)_{\varepsilon>0}$ is uniformly bounded with respect to $\varepsilon$ in $\mathcal{L}\left(L^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$: there exists $(C,N)\in\mathbb{R}_{>0}\times\mathbb{N}_{\geqslant 0}$ such that for all $a=(a_{i,j})_{1\leqslant i,j\leqslant 2}\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),$
$\left\lVert\mathrm{Op}_{\varepsilon}(a)f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant C\,\max_{1\leqslant i,j\leqslant 2}\,\max_{\left|\alpha\right|+\left|\beta\right|\leqslant N}\left\lVert\partial^{\alpha}_{x}\partial^{\beta}_{\xi}a_{i,j}\right\rVert_{\mathrm{L}^{\infty}\left(\mathbb{R}^{4}\right)}\left\lVert f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$ (7)
The Wigner transform is not a positive distribution. However, its weak limits are positive measures. This is due to the weak Gårding inequality. Adapting [11] Proposition 2.13 to the matrix case, for all $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ non negative, for all $\delta>0$, there exists $C_{\delta}>0$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
$\left\langle\mathrm{Op}_{\varepsilon}(a)f,f\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\delta+C_{\delta}\varepsilon\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
There exists a subsequence $\left(f_{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$ and a positive matrix of Radon measures $\mathrm{d}\mu=(\mathrm{d}\mu_{i,j})_{1\leqslant i,j\leqslant 2}$ such that for all $a\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
$\mathrm{I}^{\varepsilon_{n}}_{f}(a)\underset{n\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{d}\mu(x,\xi)\Big{)}.$ (8)
The measure $\mathrm{d}\mu$ is called the Wigner measure or the semiclassical measure of $\left(f^{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$. A proof of existence is recalled in Section B.1 for families valued in a separable Hilbert space.
Example 2.1.
Let us consider $\alpha\in(0,1]$, $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $x\in\mathbb{R}^{2}$,
$f^{\varepsilon}(x)\coloneqq\frac{1}{\varepsilon^{\alpha/2}}f\left(\frac{x}{\varepsilon^{\alpha}}\right),$
so ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ has only one Wigner measure that depends of $\alpha$.
If $\alpha=1/2$, then for all $(x,\xi)\in\mathbb{R}^{4}$, ${\mathrm{d}\mu(x,\xi)=\left(\int_{\mathbb{R}^{2}}f(x)\otimes\overline{f(x)}\mathrm{d}x\right)\delta_{0}(x)\otimes\delta_{0}(\xi)}$.
If $\alpha=1$, then for all $(x,\xi)\in\mathbb{R}^{4}$, ${\mathrm{d}\mu(x,\xi)=\widehat{f}(\xi)\otimes\overline{\widehat{f}(\xi)}\delta_{0}(x)\frac{\mathrm{d}\xi}{2\pi}}$.
We say a bounded family $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ is $\varepsilon$-oscillating if
$\limsup_{\varepsilon\to 0}\int_{\left|\xi\right|>R/\varepsilon}\left|\widehat{f^{\varepsilon}}(\xi)\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}\xi\underset{R\to+\infty}{\longrightarrow}0.$
We say a bounded family $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ is compact at infinity if
$\limsup_{\varepsilon\to 0}\int_{\left|x\right|>R}\left|f^{\varepsilon}(x)\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}x\underset{R\to+\infty}{\longrightarrow}0.$
If $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating and $\left(\left|f^{\varepsilon}\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}x\right)_{\varepsilon>0}$ converges weakly to $\nu$, then for all semiclassical measure $\mathrm{d}\mu$ of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$,
$\nu(x)=\mathrm{tr}\int_{\mathbb{R}^{2}}\mathrm{d}\mu(x,\mathrm{d}\xi).$
If $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity and $\left(\left|\widehat{f^{\varepsilon}}\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}\xi\right)_{\varepsilon>0}$ converges weakly to $\widetilde{\nu}$, then for all semiclassical measure $\mathrm{d}\mu$,
$\widetilde{\nu}(\xi)=\mathrm{tr}\int_{\mathbb{R}^{2}}\mathrm{d}\mu(\mathrm{d}x,\xi).$
Now, we apply these first properties on solutions to equation (1).
2.2 Application to the Dirac equation
In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. We revisit in this paragraph the notions introduced in the preceding one for this special family.
2.2.1 $\varepsilon$-oscillation of solution to Dirac equation
Lemma 2.2 (Conservation of $\varepsilon$-oscillating).
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating uniformly for $t\in[-T,T]$. In other words, uniformly for $t\in[-T,T]$,
$\limsup_{\varepsilon\to 0}\int_{\left|\xi\right|>R/\varepsilon}\left|\widehat{\psi^{\varepsilon}_{t}}(\xi)\right|^{2}\mathrm{d}\xi\underset{R\to+\infty}{\longrightarrow}0.$
Proof of Lemma 2.2.
Let $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ be $\varepsilon$-oscillating. The proof relies on the analysis of
$\psi^{\varepsilon,R}_{t}\coloneqq\chi^{R}(\varepsilon\mathrm{D}_{x})\psi_{t}^{\varepsilon},$
where ${\chi^{R}(\xi)=\chi\left(\frac{\xi}{R}\right)}$ with $\chi\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},[0,1]\right),\chi\equiv 0$ for $\left|\xi\right|<1$ and $\chi\equiv 1$ for $\left|\xi\right|>2$. We observe that $\psi^{\varepsilon,R}_{t}$ solves
$\mathrm{i}\varepsilon\partial_{t}\psi^{\varepsilon,R}_{t}=\mathrm{H}_{\varepsilon}\psi^{\varepsilon,R}_{t}+\left[\chi_{R}(\varepsilon\mathrm{D}_{x}),\begin{pmatrix}\mathrm{m}(x)&0\\ 0&-\mathrm{m}(x)\end{pmatrix}\right]\psi^{\varepsilon,R}_{t}.$
Nevertheless, the derivatives of $\mathrm{m}$ are bounded therefore there exists $C>0$ such that
$\left\lVert\frac{1}{\varepsilon}\left[\chi^{R}(\varepsilon\mathrm{D}_{x}),\begin{pmatrix}\mathrm{m}(x)&0\\ 0&-\mathrm{m}(x)\end{pmatrix}\right]\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\leqslant\frac{C}{R}.$
Then, by an energy argument, for all $t\in[-T,T]$,
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\left\lVert\psi^{\varepsilon,R}_{t}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant\frac{C}{R},$
so,
$\limsup_{\varepsilon\rightarrow 0}\left\lVert\psi^{\varepsilon,R}_{t}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $\leqslant\limsup_{\varepsilon\rightarrow 0}\left\lVert\psi^{\varepsilon,R}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\frac{C}{R}\left|T\right|,$
$\underset{R\to+\infty}{\longrightarrow}0.$
because $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\varepsilon$-oscillating. ∎
The $\varepsilon$-oscillation conservation property justifies the semiclassical approach of solutions to equation (1) for analyzing the limits of the quantities (3). Note that similar arguments give that there is also no loss of mass at infinity in configuration space in the sense that if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity, then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is compact at infinity uniformly in time $t\in[-T,T]$, for some $T>0$. The proof is in Appendix A.
As a consequence, if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is normalized, $\varepsilon$-oscillating and compact at infinity, then for all $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\left|x\right|<1$, $\chi\equiv 0$ for $\left|x\right|>2$,
$\lim_{\varepsilon\to 0}\left\langle\mathrm{Op}_{\varepsilon}\left(\chi\left(\frac{x}{R}\right)\chi\left(\frac{\varepsilon\mathrm{D}_{x}}{R}\right)\right)\psi^{\varepsilon}_{0},\psi^{\varepsilon}_{0}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\;\underset{R\to+\infty}{\longrightarrow}\;\lim_{\varepsilon\to 0}\left\lVert\psi_{0}^{\varepsilon}\right\rVert^{2}_{\mathrm{L}^{2}\left(R^{2},\mathbb{C}^{2}\right)}=1.$
So for all semiclassical measure $\mathrm{d}\mu$ of $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ normalized, $\varepsilon$-oscillating and compact at infinity,
$\mathrm{tr}\int_{\mathbb{R}^{4}}\mathrm{d}\mu(x,\xi)=1.$
2.2.2 Time-averaged Wigner measures
For taking into account the time dependence of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, we will consider time-averaged Wigner measure. We test the Wigner transform $\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}]$ against time-dependent observables, so for all $\Xi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R},\mathbb{C}\right)$,
$\int_{\mathbb{R}}\int_{\mathbb{R}^{4}}\Xi(t){\rm tr}_{\mathbb{C}^{2}}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi\mathrm{d}t\;=\;\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t.$
For the same reasons as before, we obtain the existence of a vanishing sequence $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ of positive numbers and of a measure $\mathrm{d}\mu$ on $\mathbb{R}^{5}=\mathbb{R}_{t}\times\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$ such that
$\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\;\underset{k\to+\infty}{\longrightarrow}\;\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi).$
Moreover, thanks to the control
$\left|\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi)\right|\leqslant C\ \left\lVert\Xi\right\rVert_{\mathrm{L}^{1}(\mathbb{R})}\left\lVert\mathrm{Op}_{\varepsilon}(a)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)},$
we can write $\mathrm{d}\mu(t,x,\xi)=\mathrm{d}\mu_{t}(x,\xi)\mathrm{d}t$. In other words, the measure $\mathrm{d}\mu$ is absolutely continuous in time with respect to Lebesgue measure.
2.2.3 First properties of the Wigner measure of the solution of the Dirac equation
It is known from [15, 11] that any time averaged Wigner measure, outside the curve $\mathcal{C}$, of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is completely characterized by those of the initial data. Let us first introduce some notations. Let us consider the symbol of the operator $\mathrm{H}_{\varepsilon}$ according to the Weyl quantization; for all $(x,\xi)\in\mathbb{R}^{4}$,
$\mathrm{H}(x,\xi)\coloneqq\begin{pmatrix}\mathrm{m}(x)&\xi_{1}-\mathrm{i}\xi_{2}\\ \xi_{1}+\mathrm{i}\xi_{2}&-\mathrm{m}(x)\end{pmatrix}.$
We remark that ${\det\mathrm{H}(x,\xi)=-\mathrm{m}(x)^{2}-|\xi|^{2}}$. Let $(x,\xi)\in\mathbb{R}^{4}$ such that $\det\mathrm{H}(x,\xi)\neq 0$, then $\mathrm{H}(x,\xi)$ has two different eigenvalues $\lambda_{+}(x,\xi)$ and $\lambda_{-}(x,\xi)$ defined by
[…]
Let us denote the projectors on the corresponding eigenspaces by
$\Pi_{+}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{+}(x,\xi)}\mathrm{H}(x,\xi)\quad,\quad\Pi_{-}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{-}(x,\xi)}\mathrm{H}(x,\xi).$
According to the definition of $\mathcal{C}$, we remark that
$(x,\xi)\in\mathcal{C}\iff\det\mathrm{H}(x,\xi)\neq 0.$
We recall the definition of the Poisson bracket
$\left\{f,g\right\}\coloneqq\nabla_{\xi}f\nabla_{x}g-\nabla_{x}f\nabla_{\xi}g\;,\;(f,g)\in\mathscr{C}^{1}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}\right)^{2}.$ (9)
Lemma 2.3 (Evolution of semiclassical measure outside $\mathcal{C}$).
Let $\mathrm{d}\mu_{t}$ be a semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$, then the measure $\mathrm{d}\mu_{t}$ can be decomposed outside $\mathcal{C}$ as follows : for all $(x,\xi)\in\mathbb{R}^{4}\setminus\mathcal{C}$
$\mathrm{d}\mu_{t}(x,\xi)=\Pi_{+}(x,\xi)\mathrm{d}\mu_{t}^{+}(x,\xi)+\Pi_{-}(x,\xi)\mathrm{d}\mu_{t}^{-}(x,\xi),$
where the scalar measures $\mathrm{d}\mu_{t}^{\pm}$ satisfy
$\partial_{t}\ \mathrm{d}\mu_{t}^{+}=\left\{\mathrm{d}\mu_{t}^{+},\lambda_{+}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{+}=\mathrm{tr}\Big{(}\Pi_{+}\mathrm{d}\mu_{0}\Big{)},$
$\partial_{t}\ \mathrm{d}\mu_{t}^{-}=\left\{\mathrm{d}\mu_{t}^{-},\lambda_{-}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{-}=\mathrm{tr}\Big{(}\Pi_{-}\mathrm{d}\mu_{0}\Big{)},$
and $\mathrm{d}\mu_{0}$ is the semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$.
This theorem is proved in [15, section 6]. The proof uses two major arguments. First, we have the decomposition of the measure $\mathrm{d}\mu_{t}$ as a sum $\Pi_{+}\mathrm{d}\mu_{t}^{+}+\Pi_{-}\mathrm{d}\mu_{t}^{-}$. Secondly, the equation for $\mathrm{d}\mu_{t}^{\pm}$ comes from the analysis for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}\right)$ of
$\frac{1}{\mathrm{i}\varepsilon}\Big{[}\mathrm{Op}_{\varepsilon}\left(a\Pi_{\pm}\right),\mathrm{Op}_{\varepsilon}\left(\mathrm{H}_{\varepsilon}\right)\Big{]}.$
This theorem implies that for all initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ localized around $(x_{0},\xi_{0})\notin\mathcal{C}$, the semiclassical measure evolved along the Hamiltonian trajectories associated to the functions $\lambda_{+}$ and $\lambda_{-}$. These trajectories are included in the hypersurface $\det\mathrm{H}(x,\xi)=\det\mathrm{H}(x_{0},\xi_{0})$. It means that if we localize the initial data far from the curve $\mathcal{C}$, then the semiclassical measure never approaches the curve $\mathcal{C}$ in the phase space. Our aim in the next sections is to describe the evolution of the measure $\mathrm{d}\mu_{t}$ above the curve $\mathcal{C}$.
3 Two-scale analysis of concentrations
To study the evolution of the measure $\mathrm{d}\mu_{t}$ near the curve $\mathcal{C}$, we use a two-scale analysis near $\mathcal{C}$. We use the geometric assumptions on $\mathcal{C}$ stated in the introduction to define our two-scale quantization. In geometric terms, the measure $\mathrm{d}\mu_{t}$ is a measure on the cotangent space to $\mathbb{R}^{4}$ and, as such, have the associated geometric invariance. In particular, we can operate a change of coordinates. We work locally in a tubular neighborhood of $\mathcal{C}$ where we can use normal geodesic coordinates that we introduce in Section 3.1. In these coordinates $(s,y)$, the curve $\mathrm{E}$ writes $\{y=0\}$. We use these coordinates to transform the operator and obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$. In Section 3.2 we introduce two-scale Wigner measures. Finally, we state our main results, describing these two-scale semiclassical measures for families of solutions to equation (1) in Section 3.3.
3.1 Normal geodesic coordinates
We assume assumptions 1.2 and 1.3. We will work in the normal geodesic coordinates denoted $(s,y)$ defined by (4). The curve $\mathcal{C}$ in these new coordinates is reduced to $\left\{y=\sigma=\eta=0\right\}$ where $(\sigma,\eta)$ play the role of $\xi$.
Let us define ${\mathrm{I}_{\delta_{0}}\coloneqq\left(-\frac{\delta_{0}}{2||\kappa||_{\infty}},\frac{\delta_{0}}{2||\kappa||_{\infty}}\right)}$ for $\delta_{0}\in\left(0,1\right]$ and let us consider the map
$\begin{matrix}\Phi_{\delta_{0}}:&\mathbb{R}\times\mathrm{I}_{\delta_{0}}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$
A consequence of Assumption 1.3 is that for all $\delta_{0}\in\left(0,1\right]$, the function $\Phi_{\delta_{0}}$ is a diffeomorphism from ×$\mathbb{R}\times\mathrm{I}_{\delta_{0}}$ to
$\Omega_{\delta_{0}}$$\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}_{\delta_{0}}\right\},$
=$=\left\{x\in\mathbb{R}^{2}\Big{|}\ \mathrm{d}\left(x,\mathrm{E}\right)<\frac{\delta_{0}}{2||\kappa||_{\infty}}\right\}.$
We now use $\Phi$ to obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$ as in [3].
Proposition 3.1 (Hamiltonian Normal form near the interface).
There exists a unitary operator $\mathrm{U}$ from $\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ to $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ such that $\mathrm{U}\mathrm{H}_{\varepsilon}\mathrm{U}^{-1}=\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ where the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ is acting on $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ and defined as
$\mathrm{H}_{\varepsilon}^{\mathrm{E}}\coloneqq\mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{1}-\varepsilon\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}\sigma_{3}.$
Remark.
We have the following properties.
  1. 1.
    For all $(s,y)\in\mathbb{R}\times\mathrm{I},\left|1-y\kappa(s)\right|>1/2$ therefore the previous quantities are well-defined.
  2. 2.
    The operator $\mathrm{U}$ is a matrix-valued Fourier Integral Operator. It contains the pull-back by the change of variables $\Phi$ and the multiplication by an adequate $s$-dependent matrix (a rotation and a multiplication by a $s$-phase function).
  3. 3.
    In the following, we will more generally apply the operator $\mathrm{U}$ to function of $\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ but compactly supported in $\Omega$.
Proof of Proposition 3.1.
Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$ and for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$,
$\mathrm{U}u(s,y)\coloneqq\sqrt{\frac{1-y\kappa(s)}{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\end{pmatrix}u\Big{(}\Phi(s,y)\Big{)}.$
In order to facilitate the computation, we split the operator $\mathrm{U}$ into the composition of two operators. Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$, for all $\theta\in\mathbb{R}$, for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ and for all $v\in\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$,
$\mathrm{V}_{\Phi}u(s,y)$ $\coloneqq\sqrt{1-y\kappa(s)}u\Big{(}\Phi(s,y)\Big{)},$
$\mathrm{U}_{\theta}v(s,y)$ $\coloneqq\frac{1}{\sqrt{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\end{pmatrix}v(s,y).$ (10)
The operator $\mathrm{V}_{\Phi}$ is a unitary operator from $\mathrm{L}^{2}\left(\Omega,\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ and the operator $\mathrm{U}_{\theta}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$. Notably, $\mathrm{U}=\mathrm{U}_{\theta(s)}\mathrm{V}_{\Phi}$. Moreover,
$\mathrm{V}_{\Phi}\mathrm{H}_{\varepsilon}\mathrm{V}_{\Phi}^{-1}=$ $\ \mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{3}+\frac{\varepsilon}{1-y\kappa(s)}(\gamma^{\prime}\odot\sigma)\mathrm{D}_{s}+\varepsilon(\nu\odot\sigma)\mathrm{D}_{y}$
$\ -\ \frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}(\gamma^{\prime}\odot\sigma)-\frac{\mathrm{i}\varepsilon\kappa(s)}{2(1-y\kappa(s))}(\nu\odot\sigma),$
with the notation $x\odot\sigma\coloneqq x_{1}\sigma_{1}+x_{2}\sigma_{2}$ for $x\in\mathbb{C}^{2}$ and $\sigma_{1},\sigma_{2}$ Pauli matrices. We conclude with the identities
$\begin{split}\mathrm{U}_{\theta(s)}\sigma_{3}\mathrm{U}_{\theta(s)}^{-1}=\sigma_{1},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=\sigma_{3},\\ \mathrm{U}_{\theta(s)}(\nu\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=-\sigma_{2},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\left(\mathrm{D}_{s}\mathrm{U}_{\theta(s)}^{-1}\right)=-\frac{\kappa(s)}{2}\mathrm{i}\sigma_{2}.\end{split}$
Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solves (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. Let $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\mathrm{d}(x,\mathrm{E})<\frac{1}{4||\kappa||_{\infty}}$ and $\chi\equiv 0$ for $\mathrm{d}(x,\mathrm{E})>\frac{1}{2||\kappa||_{\infty}}$, then we define for all $\varepsilon>0$,
$\varphi^{\varepsilon}_{t}\coloneqq\mathrm{U}\chi\psi^{\varepsilon}_{t}.$ (11)
Lemma 3.2 (Dirac equation in normal geodesic coordinates).
The family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$ the following equation
$\varepsilon\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi^{\varepsilon}_{t}=0.$ (12)
Proof of Lemma 3.2.
We apply $\mathrm{U}\chi$ to equation (1),
$\varepsilon\mathrm{D}_{t}\varphi_{t}^{\varepsilon}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi_{t}^{\varepsilon}=\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}.$
We compute
$\left[\mathrm{H}_{\varepsilon},\chi\right]=-\mathrm{i}\varepsilon\begin{pmatrix}0&\partial_{1}\chi-\mathrm{i}\partial_{2}\chi\\ \partial_{1}\chi+\mathrm{i}\partial_{2}\chi&0\end{pmatrix},$
where for all $j\in\left\{1,2\right\},\partial_{j}\chi\in\mathscr{C}_{c}^{\infty}(\mathbb{R}^{2},\mathbb{C})$ is supported in $\left\{x\in\mathbb{R}^{2}\ \Big{|}\ \frac{1}{4||\kappa||_{\infty}}\leqslant\mathrm{d}(x,\mathrm{E})\leqslant\frac{1}{2||\kappa||_{\infty}}\right\}$, then for all $(t,s,y)\in\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$,
$\Big{(}\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}\Big{)}(s,y)=0.$
3.2 Two-scale Wigner measures
In this section, we introduce observables that allow to concentrate on the set $\left\{y=0,\sigma=0,\eta=0\right\}$ at a precise scale, that we choose as $\sqrt{\varepsilon}$. As we will see later and as expected in view of [1], this scale is natural.
3.2.1 Two-scale symbols and their quantization
Following the course [11], we introduce
Definition (Two-scale observable).
We define the set $\mathfrak{A}$ as follows
$\mathfrak{A}\coloneqq\left\{a\in\mathscr{C}^{\infty}\Big{(}\mathbb{R}^{4}\times\mathbb{R}^{3},\mathbb{C}^{2,2}\Big{)}\ \Bigg{|}\ \begin{array}[]{c}\exists K\subset\mathbb{R}^{4}\ \text{compact},\forall z\in\mathbb{R}^{3},a(\cdot,\cdot,z)\ \text{is supported in K}\\ \exists R_{0}>0,\forall|z|>R_{0},a(\cdot,\cdot,z)=a_{\infty}\left(\cdot,\cdot,\frac{z}{|z|}\right)\end{array}\right\}.$
We can identify $\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ as a subset of $\mathfrak{A}$. We define a second scale quantization near the interface $\mathcal{C}$ for these observable by
$\begin{array}[]{ccccc}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}&:&\mathfrak{A}&\to&\mathcal{L}\big{(}\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\big{)}\\ &&a&\mapsto&\mathrm{Op}_{1}\left(a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\right)\end{array}$ (13)
In particular, we have, for all $a\in\mathfrak{A}$,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{\sqrt{\varepsilon}}\right)\right)$
$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Big{(}a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\Big{)}\mathfrak{T}_{\sqrt{\varepsilon}},\ \left(\text{semiclassical in }\sqrt{\varepsilon}\right)$ (14)
with for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $(s,y)\in\mathbb{R}^{2},$
$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}[f](s,y)=\frac{1}{\varepsilon^{1/4}}f\left(s,\frac{y}{\sqrt{\varepsilon}}\right).$ (15)
Notably, $\mathfrak{T}_{\sqrt{\varepsilon}}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Moreover, the quantization map $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}$ associates to $a$ a bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ according to the Calderón-Vaillancourt Theorem (7) applied to expression (14).
These observables allow us to differentiate three regions of the phase space $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}_{\sigma,\eta}$. Let us explain this fact. Let $a\in\mathfrak{A}$ and $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ with $\chi\equiv 1$ on $[-1/2,1/2]$. Let $\delta>0$, $R>R_{0}$ with $R_{0}$ associated to $a$ and let us consider $\varepsilon$ small enough such that
$2R\sqrt{\varepsilon}<\delta.$ (16)
Let us denote $x\coloneqq(s,y),\xi\coloneqq(\sigma,\eta)$ and $z\coloneqq\left(y,\sigma,\eta\right)$ for simplicity. We now split $a$ into three observables,
$a\left(x,\xi,z\right)=$ $\ a\left(x,\xi,z\right)\left(1-\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\right)$ $\left(\eqqcolon a^{\delta}_{\varepsilon}(x,\xi,z)\right)$ (17)
$+\ a\left(x,\xi,z\right)\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\left(1-\chi\left(\frac{\left|z\right|}{R}\right)\right)$ $\left(\eqqcolon a_{\varepsilon,\delta}^{R}(x,\xi,z)\right)$ (18)
$+\ a\left(x,\xi,z\right)\chi\left(\frac{\left|z\right|}{R}\right)$ $\left(\eqqcolon a_{\varepsilon,R}(x,\xi,z)\right).$ (19)
Let us define $\chi_{R}\coloneqq\chi\left(\frac{\left|\cdot\right|}{R}\right)$ and $\chi^{R}\coloneqq 1-\chi_{R}$. The symbol $a_{\varepsilon}^{\delta}$ through the second scale quantization is supported outside $\mathcal{C}$, at a distance $\delta$ of $\mathcal{C}$. Because of (16), $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)$ is a semiclassical operator which satisfies
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\right)\right).$
So, we can apply the semiclassical theory recalled in Section 2 to the observable $a_{\varepsilon}^{\delta}$.
In the second scale quantization, the symbol $a_{\varepsilon,R}$ is supported at distance of order $\sqrt{\varepsilon}$ of $\mathcal{C}$ and $a_{\varepsilon,\delta}^{R}$ is supported at distance larger than $\sqrt{\varepsilon}$ of $\mathcal{C}$ but smaller than $\delta$. We are particularly interested by the two regions covered by $a_{\varepsilon,\delta}^{R}$ and $a_{\varepsilon,R}$. They will describe the measure near the curve $\mathcal{C}$.
For the observable $a_{\varepsilon,\delta}^{R}$, we have
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a\left(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$ (20)
But we choose $R>R_{0}$ then
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a_{\infty}\left(s,0,0,0,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{\left|(y,\varepsilon\sigma,\varepsilon\eta)\right|}\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$
By using (14) on $a_{\varepsilon,R}$, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)$ $=\mathrm{Op}_{1}\left(a(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right),$
$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (21)
So we will consider the separable Hilbert space $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and the setting of Appendix B: the functional space $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, and the two-scale quantization defined for $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$,
$\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\coloneqq\mathrm{Op}_{1}\Big{(}b\left(s,\sqrt{\varepsilon}\sigma\right)\Big{)}.$ (22)
Let us denote, for all $(s,\sigma)\in\mathbb{R}^{2}$, the bounded operator on $\mathcal{H}$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$ by $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$. Then, for the operator $b(s,\sigma)$ defined by for all $(s,\sigma)\in\mathbb{R}^{2}$, $b(s,\sigma)\coloneqq a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (23)
In the following sections, we will take advantage of these properties of $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$.
3.2.2 Two-scale semiclassical measures
Let us first consider the stationary case. We use the observable of $\mathfrak{A}$ to analyze the oscillations at the scale $\sqrt{\varepsilon}$ and $\delta$ of uniformly bounded families $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ through $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$, respectively. Let us define for $a\in\mathfrak{A}$ the quantity
$\mathcal{I}^{\varepsilon}_{f}(a)\coloneqq\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
Theorem 3.3 (Existence of the two-scale semiclassical measure).
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
  • $\bullet$
    Two measurable map $\mathrm{d}\nu_{\infty}$ and $\mathrm{d}\gamma$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
  • $\bullet$
    A $\mathrm{d}\nu_{\infty}$-measurable map $(s,\omega)\mapsto M_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
  • $\bullet$
    A $\mathrm{d}\gamma(s,\sigma)$ measurable map $(s,\sigma)\mapsto\Gamma(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
such that for all $a\in\mathfrak{A}$ we have
$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right)$ (24)
+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$ (25)
$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$ (26)
where $\mathrm{d}\mu$ the semiclassical measure of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
The proof of this theorem is in Appendix B.
We denote by $\Gamma\mathrm{d}\gamma$ the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ with $\Gamma$ a trace one positive operator $\mathrm{d}\gamma$-measurable and $\mathrm{d}\gamma$ a non negative measure. We denote by $M_{\infty}\mathrm{d}\nu_{\infty}$ the pair $\left(M_{\infty},\mathrm{d}\nu_{\infty}\right)$ with $M_{\infty}$ a Hermitian trace one positive matrix $\mathrm{d}\nu_{\infty}$-measurable and $\mathrm{d}\nu_{\infty}$ a non negative measure.
Definition (Two-scale semiclassical measure).
We call two-scale semiclassical measures, associated with the concentration of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ near the curve $\mathcal{C}$ of order between $\sqrt{\varepsilon}$ and $\delta$ and of order $\sqrt{\varepsilon}$, respectively, the pair ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$. We will say that $\Gamma\mathrm{d}\gamma$ is a two-scale semiclassical measure at finite distance and $M_{\infty}\mathrm{d}\nu_{\infty}$ is a two-scale semiclassical measure at infinity.
Remark 3.4 (Link with the limit of the previous observable).
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3, then, with the same notations, for all $a\in\mathfrak{A}$,
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k}}^{\delta}\right)=$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right),$
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},\delta}^{R}\right)=$$\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega),$
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},R}\right)=$ $\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Example.
Let us consider, for all $(s,y)\in\mathbb{R}^{2}$,
$\phi_{\varepsilon}(s,y)=\frac{1}{\varepsilon^{1/4}}u_{\varepsilon}\left(s\right)f\left(\frac{y}{\sqrt{\varepsilon}}\right)\vec{V}$
with $\vec{V}\in\mathbb{C}^{2}$, $(f,u_{\varepsilon})\in\mathscr{S}(\mathbb{R},\mathbb{C})^{2}$ and the knowledge of $\mathrm{d}\rho$ the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$, then for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{7},\mathbb{C}^{2,2})$,
$\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)\phi_{\varepsilon},\phi_{\varepsilon}\right\rangle\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$
where
$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\frac{1}{2\pi}\Pi_{f\vec{V}}\mathrm{d}\rho(s,\sigma),$
with $\Pi_{f\vec{V}}$ the projector on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$ over the function $f\vec{V}$. Notably, if $u_{\varepsilon}$ is a wave packet, which means for all $s\in\mathbb{R}$,
$u_{\varepsilon}\left(s\right)\coloneqq\frac{1}{\varepsilon^{1/4}}u\left(\frac{s}{\sqrt{\varepsilon}}\right)$
with $u\in\mathscr{S}(\mathbb{R},\mathbb{C})$, then the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$ is $\mathrm{d}\rho(\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\mathrm{d}\sigma$ (according to the case $\alpha=1$ of Example 2.1 but with a $\sqrt{\varepsilon}$-Weyl quantization) so
$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\Pi_{f\vec{V}}\frac{\mathrm{d}\sigma}{2\pi}.$
Remark 3.5.
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3. According to the notations of Theorem 3.3, the two-scale semiclassical measures ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$ satisfy for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}a\left(s,y,\sigma,\eta\right)\mathrm{d}\mu(s,y,\sigma,\eta)\Big{)}$
+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a(s,0,0,0)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$
$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a(s,0,0,0)\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Corollary 3.6 (A condition to have $M_{\infty}\mathrm{d}\nu_{\infty}=0$).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ such that there exists $\chi$ a cut off function with $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},\mathbb{R})$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
Then, for all two-scale semiclassical measure $M_{\infty}\mathrm{d}\nu_{\infty}$ at infinity of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$,
$M_{\infty}\mathrm{d}\nu_{\infty}=0.$
Proof of Corollary 3.6.
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\chi$ a cut off function, $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
For all $a\in\mathfrak{A}$, we consider,
$\alpha(s,y,\sigma,\eta)$ $\coloneqq a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right),$
$\beta(y,\sigma,\eta)$ $\coloneqq 1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right).$
According to pseudo-differential calculus and of the inequality (16), we fix $\delta$ and let $\varepsilon$ goes to 00 then $R$ goes to +$\infty$+ ∞, so
$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\beta(y,\sigma,\eta)\Big{)}=$ $\ \mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}$
$+\ \mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\sqrt{\varepsilon}\right)+\mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\frac{1}{R}\right).$
Moreover,
$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}.$
By application of the Calderón-Vaillancourt Theorem (7),
$\left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$ $=\left\lVert\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$
$\leqslant C(a)$
where $C$ does not depend of $\varepsilon$. Then
$\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=$ $\ \Big{\langle}\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon},\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\Big{\rangle}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
$\left|\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)\right|\leqslant$ $\ \left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
$\leqslant$ $\ C(a)\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right).$
Then,
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=0.$
We conclude by using Remark 3.4. ∎
3.2.3 Time dependent two-scale semiclassical measures
As in Section 2.2.2, one can extend the definition of two-scale semiclassical measure to time dependent families. More precisely, we consider families ${\left(f^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ that are uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$. Let $a\in\mathfrak{A},\Xi\in\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$, we define the quantity
$\mathcal{I}^{\varepsilon}_{f}(\Xi,a)\coloneqq\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f_{t}^{\varepsilon},f_{t}^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}\mathrm{d}t.$
As before, we are interested by the limit of $\mathcal{I}^{\varepsilon}_{f}(\Xi,a)$ when $\varepsilon$ goes to 00. Moreover, if $a$ does not depend of $z$, then the limit will be the semiclassical measure as defined in (8). Adapting the context of  [10] to time dependent families, one defines two-scale semiclassical measures over the curve $\mathcal{C}$. Indeed, by the same argument of Section 2.2.2, we can extend Theorem 3.3 to time families. Therefore, such as before, the limit measures will be absolutely continuous in time variable with respect to Lebesgue measure.
Theorem 3.7 (Existence of the time-averaged two-scale semiclassical measure).
Let ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and
  • $\bullet$
    Two measurables map $t\mapsto\mathrm{d}\nu^{t}_{\infty}$ and $t\mapsto\mathrm{d}\gamma^{t}$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
  • $\bullet$
    A $\mathrm{d}\nu^{t}_{\infty}\otimes\mathrm{d}t$-measurable map $(t,s,\omega)\mapsto M^{t}_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
  • $\bullet$
    A $\mathrm{d}\gamma^{t}(s,\sigma)\otimes\mathrm{d}t$ measurable map $(t,s,\sigma)\mapsto\Gamma^{t}(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
such that for all $(a,\Xi)\in\mathfrak{A}\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{f}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu_{t}(s,y,\sigma,\eta)\right)\mathrm{d}t$
+×$+\ \int_{\mathbb{R}}\int_{\mathbb{R}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$+\ \int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t,$
where $\mathrm{d}\mu_{t}$ is the semiclassical measure of ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
Remark (Geometrical point of view).
In the parametrization $(s,y)$ of ×$\mathbb{R}\times\mathrm{I}$, the variables $(s,y,\sigma,\eta)$ parametrized $\mathrm{T}^{*}\mathbb{R}^{2}$. Moreover the variable $s$ parametrizes the curve $\mathrm{E}\subset\mathbb{R}^{2}$. The variable $y$ parametrizes the normal fiber to the curve $\mathrm{E}$ at the point $\gamma(s)$. In the computation of $\mathrm{Tr}$, the space $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ can be understand as the space $\mathrm{L}^{2}(N_{\gamma(s)}\mathrm{E},\mathbb{C}^{2})$. The variable $\sigma$ denotes the variable of the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\mathrm{E}$. For $\gamma(s)\in\mathrm{E}$ and $\mathrm{T}_{s}\mathrm{E}$, the tangent to $\mathrm{E}$ above $s$, the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\gamma(s)$ is $(\mathrm{T}_{s}\mathrm{E})^{*}\sim\mathbb{R}$. The measure $\mathrm{d}\gamma^{t}$ defines a measure on $\mathrm{T}^{*}\mathrm{E}$.
Since $\mathcal{C}=\{y=\eta=\sigma=0\}\subset\mathrm{T}^{*}\mathbb{R}^{2}$, $\mathcal{C}\times\mathbb{S}^{2}$ coincides to the normal bundle in sphere to $\mathcal{C}$: $\mathrm{S}\mathcal{C}$. The fibers of $\mathrm{S}\mathcal{C}$ over $\rho\in\mathcal{C}$ is obtained as follows : let $\mathrm{T}_{\rho}\mathcal{C}$ be the tangent space to $\mathcal{C}$ over $\rho$, define $\mathrm{N}_{\rho}\mathcal{C}$ the subspace of $\left(\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)\right)^{*}$ consisting of the linear forms on $\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)$ which vanish on $\mathrm{T}_{\rho}\mathcal{C}$, $S_{\rho}\mathcal{C}$ is obtained by taking the quotient of $\mathrm{N}_{\rho}\mathcal{C}$ by the action of $\mathbb{R}_{\geqslant 0}$ by homotheties.
3.3 Main result
Now we will describe the semiclassical measure of Theorem 3.7 for the solution to equation (1). Let us define some quantities involved in the main theorem. Let us define the Hermite functions $\left(\mathfrak{h}_{m}\right)_{m\in\mathbb{N}_{\geqslant 0}}$ as the Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}\right)$ satisfying for all $y\in\mathbb{R}$ and for all $m\in\mathbb{N}_{\geqslant 0}$,
$-\frac{\mathrm{d}^{2}\mathfrak{h}_{m}}{\mathrm{d}y^{2}}(y)+y^{2}\mathfrak{h}_{m}(y)=(2m+1)\mathfrak{h}_{m}(y).$ (27)
Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$,
$g_{0}^{s}(y)\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix}\;\;$ $,\;\;g_{n}^{s,\sigma}(y)\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}}\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(s,\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},$ (28)
$\Pi_{0}(s)\coloneqq g_{0}^{s}\otimes g_{0}^{s}\;\;$ $,\;\;\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$ (29)
where $\left(\lambda_{n}\right)_{n\in\mathbb{Z}}$ were defined in (5), $r(s)\coloneqq|\nabla\mathrm{m}(\gamma(s))|^{1/2}$. Although the functions $g_{0}^{s}$ and $\Pi_{0}$ do not depend on $\sigma$, we also will denote them by $g_{0}^{s,\sigma}$ and $\Pi_{0}(s,\sigma)$.
Let us consider for all $s\in\mathbb{R}$ and for all $z\in\mathbb{R}^{3}$,
$\Pi_{\pm}^{\infty}(s,z)\coloneqq\frac{1}{2}\mathrm{Id}\pm\frac{1}{2\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}}}\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}$ (30)
We are going to use the two-scale semiclassical measures of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to describe the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ near the curve $\mathcal{C}$. We already know its semiclassical measure and its evolution outside $\mathcal{C}$. Therefore, we will focus on terms (25) and (26).
Theorem 3.8 (Evolution of the two-scale semiclassical measure in normal geodesic coordinates).
Assume Assumption 1.2 and 1.3. Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)\right)$ defined in (11). Let ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ be a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, then we have the following properties.
  1. 1.
    There exists a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ such that
    $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}.$ (31)
    Moreover, for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
    $\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
    with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
  2. 2.
    There exists a family of non negative scalar Radon measure $\mathrm{d}\nu^{t}_{\pm}$ such that
    $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}=\Pi_{+}^{\infty}\mathrm{d}\nu^{t}_{+}+\Pi_{-}^{\infty}\mathrm{d}\nu^{t}_{-}.$ (32)
    Moreover, $\mathrm{d}\nu_{\pm}^{t}$ satisfies the invariance $\mathrm{div}_{\omega}\left(\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu_{\pm}^{t}(s,\omega)\right)=0,$ with for all $s\in\mathbb{R}$ and for all $\omega\coloneqq\left(\omega_{y},\omega_{\sigma},\omega_{\eta}\right)\in\mathbb{S}^{2}$,
    $\vec{V}^{\infty}(s,\omega)\coloneqq\begin{pmatrix}\left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
The proof of this theorem is in Section 4.3.
Remark (Continuity of the two-scale semiclassical measure at finite distance).
From the evolution (6), we can deduce that the measure $t\mapsto\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous and entirely determined by the evolution equation and the initial data.
Remark (Value at zero of the two-scale semiclassical measure at finite distance).
The value at zero of $\Gamma^{t}\mathrm{d}\gamma^{t}$ is not necessarily the two-scale semiclassical measure at finite distance of the initial condition. As we will see in the proof, $\Gamma^{t}(s,\sigma)$ commutes with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ because of the PDE satisfied by $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$. Nevertheless, there exists initial condition such that its two-scale semiclassical measure at finite distance does not commute with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$. We will see an example in Remark 4.4.
Corollary 3.9 (Evolution consequence).
With the notation of Theorem 3.8, if
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1,$
with $\mathrm{d}\mu_{0}$ a semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to the same vanishing sequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of the two-scale semiclassical measure $\left(M^{0}_{\infty}\mathrm{d}\nu^{0}_{\infty},\Gamma^{0}\mathrm{d}\gamma^{0}\right)$, then, for almost all $t$,
$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
Proof of Corollary 3.9.
With the notation of Theorem 3.8, the evolution of $(\mathrm{d}\mu_{t},\Gamma^{t}\mathrm{d}\gamma^{t})$ implies that for all $t\in\mathbb{R}$,
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}$ $=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)},$
$\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)$ $=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma).$
By taking $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}\in\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ normalized such that
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1=\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{2}^{2},$
we have the following equality because of the preservation of the $\mathrm{L}^{2}$-norm
$\left\lVert\psi^{\varepsilon}_{t}\right\rVert_{2}^{2}=$ $\ \int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}+\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}M^{t}_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)$
$+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma),$
$=$ $\ 1$
Therefore, for almost all $t$,
$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
The evolution of the two-scale semiclassical measure in normal geodesic coordinates allows us to describe the general evolution of any initial normalized condition for equation (1).
Theorem 3.10 (Evolution of the semiclassical measure of solution to equation (1)).
Assume Assumption 1.2 and 1.3. Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solves (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
  • $\bullet$
    a semiclassical measure $\mathrm{d}\mu_{t}$, evolving as described in Lemma 2.3,
  • a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ satisfying (6),
  • $\bullet$
    a two-scale semiclassical measure ${M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}}$ at infinity described in Theorem 3.8,
such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\ \int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\Big{(}a(x,\xi)\mathrm{d}\mu_{t}(x,\xi)\Big{)}\mathrm{d}t+\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s),0\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t,$
with for all $s\in\mathbb{R}$,
$\mathrm{d}\rho_{t}(s)=\int_{\mathbb{S}^{2}}\mathrm{U}_{\theta(s)}^{\dagger}M^{t}_{\infty}(s,\omega)\mathrm{U}_{\theta(s)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)+\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
where $\mathrm{U}_{\theta(s)}$ is defined in (10).
The proof of this theorem is in Section 4.3.
4 Proofs of the main result
4.1 Change of variable on semiclassical measure
First of all, we will see the consequences of the change of variable from Proposition 3.1 on the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ solution to equation (1) and the link with the two-scale semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ defined in (11).
Lemma 4.1 (Semiclassical measure and normal geodesic coordinates).
All semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ solution to equation (1) associated to $(\varepsilon_{k})_{k\in\mathbb{N}}$ satisfies
$\mathrm{d}\mu_{t}\mathrm{d}t\mathds{1}_{\mathcal{C}}=\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}\mathrm{d}t\mathds{1}_{\mathcal{C}}$
where $\mathrm{U}_{\theta(s)}$ is defined in (10) and $\mathrm{d}\rho_{t}$ is a semiclassical measure on $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}_{\sigma,\eta}^{2}\times\mathbb{R}_{t}$ associated to $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$.
Proof of Lemma 4.1.
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ such that $\mathrm{supp}(a)\subset\Omega_{1/2}\times\mathbb{R}^{2}$ and $\chi$ the cut-off function taken in (11), then for all $N\in\mathbb{N}$,
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $=\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\chi\psi^{\varepsilon}_{t},\chi\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\Omega_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right)$
$=\left\langle\mathrm{U}\mathrm{Op}_{\varepsilon}\left(a\right)\mathrm{U}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right).$
Since $\rm U=U_{\theta(s)}V_{\Phi}$, $\rm U_{\theta(s)}$ being the multiplication by a function depending only of the variable $s$ and $\rm V_{\Phi}$ a change of variables, standard results on semiclassical pseudodifferential calculus ([27, Theorem 9.3]) imply
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle=$ $\ \left\langle\mathrm{U}_{\theta(s)}\mathrm{Op}_{\varepsilon}\Big{(}\mathrm{V}_{\Phi}a\mathrm{V}_{\Phi}^{\dagger}\Big{)}\mathrm{U}_{\theta(s)}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon\right)$
$=$ $\ \left\langle\mathrm{Op}_{\varepsilon}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}$
$+\mathcal{O}\left(\varepsilon\right).$
For simplicity, we assume $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$. Passing to the limit, we obtain
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\right)$
=×$=\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\mathrm{U}_{\theta(s)}\Big{)}\right)$
$=\int_{\Omega_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a(x,\xi)\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}(x,\xi)\right).$
We conclude by taking the restriction of the semiclassical measure $\mathrm{d}\rho_{t}$ to the curve $\mathcal{C}$. ∎
According to the definition (15) of the rescaling operator $\mathfrak{T}_{\sqrt{\varepsilon}}$, we define the family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ by
$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\left[u^{\varepsilon}_{t}\right](s,y)=\varphi_{t}^{\varepsilon}(s,y)\quad,\quad\forall(t,s,y)\in\mathbb{R}\times\mathbb{R}\times\mathrm{I}_{1/2}.$
The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is defined on ×$\mathbb{R}\times\mathrm{I}_{\frac{1}{2\sqrt{\varepsilon}}}$ but we will study it on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$ with
$\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}\coloneqq\left(-\frac{\delta}{\sqrt{\varepsilon}},\frac{\delta}{\sqrt{\varepsilon}}\right)$
where we consider $\delta\leqslant\frac{1}{4\left|\left|\kappa\right|\right|_{\infty}}$.
Lemma 4.2 (Double scaling operator).
The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\begin{split}\left(\sqrt{\varepsilon}\mathrm{D}_{t}+y\left|\nabla\mathrm{m}(\gamma(s))\right|\sigma_{1}-\mathrm{D}_{y}\sigma_{2}+\frac{\sqrt{\varepsilon}}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\varepsilon y\frac{\mathrm{i}\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right)u_{t}^{\varepsilon}=-\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}u_{t}^{\varepsilon},\end{split}$ (33)
where for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta},$
$\begin{split}\big{(}\mathfrak{R}_{2}^{\varepsilon}u_{t}^{\varepsilon}\big{)}(s,y)\coloneqq&y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}u_{t}^{\varepsilon}(s,y).\end{split}$
Proof of Lemma 4.2.
On the one hand, we rewrite the equation (12) with the definition of $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\left[\varepsilon\mathrm{D}_{t}+\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)\sigma_{1}-\sqrt{\varepsilon}\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon^{3/2}y\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right]u^{\varepsilon}_{t}=0.$
On the other hand, for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)=$ $\ \underbrace{\mathrm{m}(\gamma(s))}_{=0}+\sqrt{\varepsilon}y\underbrace{\nabla\mathrm{m}(\gamma(s))\cdot\nu(s)}_{=|\nabla\mathrm{m}(\gamma(s))|}$
$+\ \varepsilon y^{2}\int_{0}^{1}(1-\tau)\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau.$
The operator $\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}$ will play a role of rest in the determination of $\Gamma\mathrm{d}\gamma$, the two-scale semiclassical measure at finite distance.
For now, we do some remarks about the two-scale operator. Indeed, we remark that $\sqrt{\varepsilon}$ is in front of $\mathrm{D}_{s}$ but there is no $\sqrt{\varepsilon}$ in front of $\mathrm{D}_{y}$. It fits with the two-scale operators that we have introduced in Section 3.2. It is a semiclassical operator for the scale $\sqrt{\varepsilon}$ in the variable $s$, without scale in $y$. Because of the second scale of the problem, we will consider the symbol of the principal term of equation (33) according only to the first variable $s$, keeping the operator character over the variable $y$. Let us define, for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\coloneqq\begin{pmatrix}\sigma&y|\nabla\mathrm{m}(\gamma(s))|+\mathrm{i}\mathrm{D}_{y}\\ y|\nabla\mathrm{m}(\gamma(s))|-\mathrm{i}\mathrm{D}_{y}&-\sigma\end{pmatrix},$ (34)
the symbol of an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ as in the end of the Section 3.2.1 in (22). From a spectral analysis (Appendix C), the eigenvalues of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ were defined in (5) with the associated normalized eigenvectors (28) and the associated eigenprojectors (29).
We also define the symbol of $\mathfrak{R}^{\varepsilon}_{2}$ for the two-scale quantization (22) as the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{\varepsilon}_{2}(s,\sigma)\coloneqq y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}.$
Moreover, we define the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{\varepsilon}_{3}(s,\sigma)\coloneqq\frac{y\kappa(s)\sigma}{1-\sqrt{\varepsilon}y\kappa(s)}\sigma_{3}.$
We define $\mathfrak{r}^{\varepsilon}\coloneqq\mathfrak{r}^{\varepsilon}_{2}+\mathfrak{r}^{\varepsilon}_{3}$. With the notation (22), the equation (33) writes
$\sqrt{\varepsilon}\mathrm{D}_{t}u^{\varepsilon}_{t}+\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}+\sqrt{\varepsilon}\mathfrak{r}^{\varepsilon}\Big{)}u^{\varepsilon}_{t}=0.$
4.2 The two-scale semiclassical measures in normal geodesic coordinates
In this section we will describe the evolution of the two-scale semiclassical measures defined in Theorem 3.7 for $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (12). In fact, in Section 4.2.1, we describe the evolution of the measure at finite distance $\Gamma^{t}\mathrm{d}\gamma^{t}$ and in Section 4.2.2, we describe the evolution of the measure at infinity $M_{\infty}^{t}\mathrm{d}\nu_{\infty}^{t}$. In Section 4.3, we prove Theorems 3.8 and 3.10. In Section 4.4, we prove Theorem 1.4.
4.2.1 Two-scale Wigner measure at finite distance
We prove here the part one of Theorem 3.8. We recall the following lemma of [16].
Lemma 4.3 (Poisson bracket and projector properties).
Let $\left(\Pi_{n}\right)_{n\in\mathbb{N}}$ be a family of orthogonal projectors such that for all $(n,k)\in\mathbb{N}^{2},\Pi_{n}\Pi_{k}=\delta_{k,n}$. Then for all $(j,k,\ell)\in\mathbb{N}^{3}$
$\Pi_{k}\Big{(}\left\{\Pi_{j},\Pi_{\ell}\right\}-\left\{\Pi_{\ell},\Pi_{j}\right\}\Big{)}\Pi_{k}=0$
Proof of part one of Theorem 3.8.
Assume Assumption 1.2 and 1.3. For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $\Gamma^{t}\mathrm{d}\gamma^{t}$. Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2. We consider the part (19) of $a$ denoted $a_{\varepsilon,R}$. Let us consider $b(s,\sigma)\coloneqq a(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, the operator acting on $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and obtained by Weyl quantization. The link between the second scale quantization of $a_{\varepsilon,R}$ (13) and the two-scale quantization of $b$ (22) was established in (23). By considering $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (33) as a family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$, we have
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}\Big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}\mathfrak{r}^{\varepsilon}\big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
Because of (9), we have
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\left(\frac{1}{2\mathrm{i}}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\frac{1}{2\mathrm{i}}\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}+\left[b,\mathfrak{r}^{\varepsilon}\right]\right)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
  • $\bullet$
    Step 1 : Structure of $\Gamma^{t}$
Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=$ $\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
$+\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathfrak{r}^{\varepsilon}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right)$
$=$ $-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right).$
But $b$ is compactly supported in $\left\{\left|(y,\sigma)\right|\leqslant R\right\}$ and for all $(s,\sigma,y)\in\mathbb{R}\times\left(-R,R\right)^{2}$,
$y\underbrace{\frac{\kappa(s)}{1-\sqrt{\varepsilon}y\kappa(s)}}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}\sigma=\mathcal{O}\left(R^{2}\right),\quad\quad y^{2}\underbrace{\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}=\mathcal{O}\left(R^{2}\right).$
Then
$\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b,\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
By passing to the limit, we have, for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
00 0 $=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,\mathrm{D}_{y}),\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t$
$=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}\Bigg{(}a(s,0,0,y,\sigma,\mathrm{D}_{y})\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t.$
So, $t$ a.e. $\mathrm{d}\gamma^{t}$ a.e. $\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}=0$, which implies (31).
  • $\bullet$
    Step 2 : Equation of $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$
By Step $1$, for all $(k,\ell)\in\mathbb{Z}^{2}$, with $k\neq\ell$ and for all $\beta$ function from $\mathbb{R}_{s}\times\mathbb{R}_{\sigma}$ to $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ such that $\Pi_{k}\beta\Pi_{\ell}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, we have
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Pi_{k}\beta\Pi_{\ell}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=0.$
Then we can consider symbol of the form $\Pi_{\ell}\beta\Pi_{\ell}$. Since the projector $\Pi_{\ell}$ has rank one, we have
$\Pi_{\ell}\beta\Pi_{\ell}=\mathrm{Tr}\Big{(}\Pi_{\ell}\beta\Pi_{\ell}\Big{)}\Pi_{\ell},$
where $\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\big{(}\Pi_{\ell}\beta\Pi_{\ell}\big{)}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Therefore we will consider symbols of the form $\beta\Pi_{\ell}$ with $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Let $n\in\mathbb{Z}$ and $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Let us define for all $(s,\sigma)\in\mathbb{R}^{2}$,
$b_{n}(s,\sigma)\coloneqq\beta(s,\sigma)\Pi_{n}(s,\sigma).$
The form of $b_{n}$ implies $[b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}]=0$. Then
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
We first show that the term with $\mathfrak{r}^{\varepsilon}$ will vanish as $\varepsilon$ goes to 00. By definition of $\Gamma^{t}\mathrm{d}\gamma^{t}$, we have
$\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,D_{y}),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)$
$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}\beta(s,\sigma)\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\beta(s,\sigma)\mathrm{Tr}\Bigg{(}\Big{[}\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
where the second equality holds because of Step $1$ and the chosen shape of $a$ and the last equality holds because $\beta$ is scalar and $\mathfrak{r}^{0}$ is the value at zero of $\mathfrak{r}^{\varepsilon}$ given by for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{0}(s,\sigma)=\mathfrak{r}^{0}_{2}(s,\sigma)+\mathfrak{r}^{0}_{3}(s,\sigma)=y^{2}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,0)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\sigma_{1}+y\kappa(s)\sigma\sigma_{3}.$
Then, we recall, for all $m\in\mathbb{Z}$,
$\mathrm{Tr}\Big{(}\big{[}\Pi_{n},\mathfrak{r}^{0}\big{]}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\Pi_{n}\mathfrak{r}^{0}\Pi_{m}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{m}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}\\ =\delta_{n=m}\left(\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}\right)=0.\$
Let $k\in\mathbb{Z}$, let us compute $\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}$,
$\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}=$ $\ \Pi_{k}\left\{\sum_{m\in\mathbb{Z}}\lambda_{m}\Pi_{m}\,,\,\beta\Pi_{n}\right\}\Pi_{k}$ (35)
$=$ $\ \Pi_{k}\sum_{m\in\mathbb{Z}}\Bigg{(}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}+\lambda_{m}\left\{\Pi_{m},\beta\right\}\Pi_{n}+\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Bigg{)}\Pi_{k}$
$\ +\Pi_{k}\sum_{m\in\mathbb{Z}}\Big{(}\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Big{)}\Pi_{k}.$
Let $m\in\mathbb{Z}$,
$\Pi_{k}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}\Pi_{k}$ $=\mathds{1}_{k=m=n}\left\{\lambda_{n},\beta\right\}\Pi_{n},$
[…]
$\Pi_{k}\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k}$ $=\lambda_{m}\beta\Pi_{k}\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k},$
[…]
where the two vanishing equations hold because $\lambda_{n}$ and $\beta$ are scalar and $\Pi_{k}\partial\Pi_{m}\Pi_{k}$ is always equal to zero. Moreover, according to Lemma 4.3,
$\lambda_{m}\beta\Pi_{k}\Big{(}\left\{\Pi_{m},\Pi_{n}\right\}-\left\{\Pi_{n},\Pi_{m}\right\}\Big{)}\Pi_{k}=0\,,\,\text{for all}\ m\in\mathbb{Z}.$
Therefore,
$\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}=\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}.$
Then,
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=\frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\ +\ \mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
By passing to the limit,
$\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi^{\prime}(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Big{)}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\left\{\lambda_{n},\beta\right\}(s,\sigma)\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t.\\ =-\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\beta(s,\sigma)\left\{\lambda_{n},\mathrm{d}\gamma^{t}_{n}\right\}(s,\sigma)\mathrm{d}t.\$
Then $\mathrm{d}\gamma_{n}^{t}$ satisfies the equation (6) and is notably continuous on $\mathbb{R}$. It remains to prove that the value at zero of $\mathrm{d}\gamma_{n}^{t}$ coincides with the measure of the initial condition $\mathrm{d}\gamma_{n,0}$. Let $T\in\mathbb{R}$, on the one hand,
$\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{T},u^{\varepsilon}_{T}\right\rangle-\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{0},u^{\varepsilon}_{0}\right\rangle.$
On the other hand,
$\left|\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t\right|$ $\leqslant\frac{1}{2}\int_{0}^{T}\left|\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\right|\mathrm{d}t\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right),$
$\leqslant\frac{C\left(\left\{\lambda_{n},\beta\right\}\Pi_{n}\right)T}{2}\left\lVert u^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}}^{2}\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right).$
where the last inequality holds because of the Calderón-Vaillancourt Theorem (7) and the preservation of the $\mathrm{L}^{2}$-norm of the solution. Then, by passing to the limit, we have
$\begin{split}\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{t}^{\varepsilon}}\left(a_{\varepsilon,R}\right)&=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma_{n}^{0}(s,\sigma),\\ &=\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{0}^{\varepsilon}}\left(a_{\varepsilon,R}\right).\end{split}$
Therefore, for all $n\in\mathbb{Z},\mathrm{d}\gamma_{n}^{t}$ is continuous and the value at zero coincides with the initial condition. Moreover, the equality $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}$ holds as trace class operator so, thanks to Lebesgue dominated convergence Theorem, $\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous on $\mathbb{R}$. ∎
Remark 4.4 (Zero value of the two-scale semiclassical measure at finite distance).
We can construct the initial data with inspiration of the proof of the Lemma B.1 where we construct semiclassical measures for families valued in a separable Hilbert space.
Let $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ be a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathbb{C}\right)$ with the knowledge of $\mathrm{d}\rho$, the semiclassical measure of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$. The function $(s,\sigma)\mapsto g_{1}^{s,\sigma}$ is smooth, bounded and with all its derivatives bounded uniformly in $y$, therefore we can consider its Weyl $\sqrt{\varepsilon}$-quantization according to the variable $s$. Let us consider initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ such that for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{1/2}$,
$u_{0}^{\varepsilon}(s,y)=g_{0}^{s}(y)f(s)+\Big{(}\mathrm{Op}_{\sqrt{\varepsilon}}^{(s)}\big{(}g_{1}^{s,\sigma}(y)\big{)}f\Big{)}(s),$
where the Weyl $\sqrt{\varepsilon}$-quantization of $g_{1}^{s,\sigma}$ is taken only in the $s$ variable. Then the two-scale semiclassical measure at finite distance of this initial data is for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\frac{1}{2\pi}\Big{(}\Pi_{0}(s)+\Pi_{1}(s,\sigma)+\Pi_{1,0}(s)+\Pi_{0,1}(s,\sigma)\Big{)}\mathrm{d}\rho(s,\sigma),$
where for all $(k,\ell)\in\left\{0,1\right\}^{2}$ and for all $(s,\sigma)\in\mathbb{R}^{2},\Pi_{k,\ell}(s,\sigma)=g^{s,\sigma}_{k}\otimes g^{s,\sigma}_{\ell}$. Notably, it does not commute with $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$.
4.2.2 Two-scale Wigner measure at infinity
We prove here the part two of Theorem 3.8. We consider the symbol of $\mathrm{H}^{\mathrm{E}}_{\varepsilon}$ according to the second scale quantization (13), given by for all $(s,y)\in\mathbb{R}^{2}$ and $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
$\mathrm{T}^{\mathrm{E}}(s,y,z)\coloneqq z_{y}\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\sigma_{1}-z_{\eta}\sigma_{2}+\frac{z_{\sigma}}{1-y\kappa(s)}\sigma_{3}$
which has two eigenvalues :
$\lambda_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\pm\sqrt{\frac{1}{(1-y\kappa(s))^{2}}z_{\sigma}^{2}+z_{y}^{2}\left|\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\right|^{2}+z_{\eta}^{2}}$
which are homogeneous functions of order 1 in $z$. The eigenprojector $\Pi_{\pm}^{\mathrm{E}}(s,y,z)$ defined by
$\Pi_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\frac{1}{2}\mathrm{Id}+\frac{1}{2\lambda_{\pm}^{\mathrm{E}}(s,y,z)}\mathrm{T}^{\mathrm{E}}(s,y,z)$
are homogeneous functions of order 0 in $z$.
Moreover, we denote the principal symbol of the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ associated to the second scale quantization (13) by for all $(s,y)\in\mathbb{R}^{2}$ and for all $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
$\mathcal{T}^{\infty}_{\mathrm{E}}(s,z)\coloneqq\mathrm{T}^{\mathrm{E}}(s,0,z)=\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}.$
Its eigenvalues are
$\lambda_{\pm}^{\infty}(s,z)\coloneqq\pm\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}},$
and the associated eigenprojectors $\Pi_{\pm}^{\infty}(s,z)$ were defined in (30).
Then, the equation (12) writes on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$
$\sqrt{\varepsilon}\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\Big{(}\mathrm{T}^{\mathrm{E}}\Big{)}\varphi^{\varepsilon}_{t}=0.$
Proof of part two of Theorem 3.8.
Assume Assumption 1.2 and 1.3. For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}$. Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2. We consider the part (18) of $a$ divided by $\lambda^{\infty}_{+}$. The symbol $\mathrm{T}^{\mathrm{E}}$ is homogeneous in $z$ of order $1$ and the symbol $a_{\varepsilon,\delta}^{R}/\lambda^{\infty}_{+}$ is homogeneous in $z$ of order $-1$. Because of (18), we consider symbols on a subset of $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant\frac{2\delta}{\sqrt{\varepsilon}}\right\}$. Moreover, we have the following properties.
  1. 1.
    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}$ is homogeneous in $z$ of order $-2$.
  2. 2.
    $\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}$ is homogeneous in $z$ of order $-1$.
  3. 3.
    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}$ is homogeneous in $z$ of order 00.
Then, according to standard semiclassical calculus,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)$ $=\mathcal{O}\left(\frac{1}{R^{2}}\right),$
$\sqrt{\varepsilon}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)$ $=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right),$
$\varepsilon\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{(s,y),(\sigma,\eta)}\right)$ $=\mathcal{O}_{R}\left(\varepsilon\right).$
So
$\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)$
$+\frac{1}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}\right)+\mathcal{O}\left(\frac{1}{R^{4}}\right)$
$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)+\mathcal{O}_{R}\left(\varepsilon\right).$
Considering $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (12), we have
$\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=\left\langle\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle,\\ =\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\\ +\mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right).\$
  • $\bullet$
    Step 1 : Structure of $M_{\infty}^{t}$
Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t\\ +\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\ \mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right)\\ =-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t.\$
But $\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]=\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\varepsilon,\delta}^{R}$ and $\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]\in\mathfrak{A}$, then by passing to the limit, we deduce that for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a_{\infty}(s,0,0,0,\omega),\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{T}^{\mathrm{E}}(s,0,\omega)\right]M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{tr}\Bigg{(}a_{\infty}(s,0,0,0,\omega)\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t.$
So, for almost all $t\in\mathbb{R}$, we have $\mathrm{d}\nu_{\infty}^{t}\ \mathrm{a.e.}$ $\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}=0$ which implies (32).
  • $\bullet$
    Step 2 : Equation of $\mathrm{d}\nu_{\infty}^{t}$.
Let $\beta\in\mathfrak{A}$ and $a$ defined by for all $(s,y,\sigma,\eta,z)\in\mathbb{R}^{7}$,
$a(s,y,\sigma,\eta,z)\coloneqq\lambda^{\infty}_{\pm}(s,z)\beta(s,y,\sigma,\eta,z)\Pi_{\pm}^{\mathrm{E}}(s,y,z).$
The form of $a$ implies $\left[a,\mathrm{T}^{\mathrm{E}}\right]=0$. Then
$\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}\left(\frac{1}{R^{2}}\right)$
$+\frac{\sqrt{\varepsilon}}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{s,z_{\sigma}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\varepsilon\right).$
Because of the homogeneity in $z$, we have $\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\in\mathfrak{A}$. Moreover,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}\right)=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left(\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\right)_{\varepsilon,\delta}^{R}\right)+\frac{1}{R}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)$
$+\frac{\sqrt{\varepsilon}}{\delta}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right).$
The function $\partial\chi^{R}$ is compactly supported in the ring $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant R\right\}$ and the function $a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}$ belongs to $\mathfrak{A}$ so
$\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)_{\infty}=0.$
The operator $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)$ is uniformly bounded in $\varepsilon$ then
$\lim_{\varepsilon\to 0}\frac{\sqrt{\varepsilon}}{\delta}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)\varphi_{t}^{\varepsilon},\varphi_{t}^{\varepsilon}\right\rangle=0.$
As in the computation of (35), we consider the quantity $\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}$ then
$\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}=\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}.$
Then, by passing to the limit, we have
00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left(\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}\right)_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t,$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\Pi_{+}^{\infty}(s,\omega)\Bigg{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
but for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
$\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{y}}(s,\omega)=\omega_{y}r(s)^{4},\quad\quad\quad\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{\eta}}(s,\omega)=\omega_{\eta}.$
Then
$\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\Big{(}\vec{V}_{+}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t=0,$
with
$\vec{V}_{+}(s,\omega)=\begin{pmatrix}-\omega_{\eta}\\ 0\\ \omega_{y}r(s)^{4}\end{pmatrix}$
and $\beta_{\infty}$ a function on the sphere with $\nabla_{\omega}\beta_{\infty}$ is a vector field over the sphere. Therefore we can rewrite $\vec{V}_{\pm}\cdot\nabla_{\omega}\beta=\vec{V}_{\pm}^{\infty}\cdot\nabla_{\omega}\beta$ where $V_{\pm}^{\infty}$ is the component of $\vec{V}_{\pm}$ which is tangential to the sphere. The vector $\vec{V}_{\pm}^{\infty}$ defines a vector field over the sphere and is given by for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
$\vec{V}^{\infty}(s,\omega)\coloneqq\vec{V}_{+}(s,\omega)-\Big{(}\vec{V}_{+}(s,\omega)\cdot\omega\Big{)}\omega=\begin{pmatrix}\left(\left(1-r(s)^{4}\right)\omega_{y}^{2}-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\left(1-r(s)^{4}\right)\omega_{\eta}^{2}+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
Because of the equality $\lambda^{\infty}_{+}=-\lambda^{\infty}_{-}$, we have the same invariance for $\mathrm{d}\nu_{-}^{t}$. So, $t$ almost everywhere,
×$\int_{\mathbb{R}\times\mathbb{S}^{2}}\Big{(}\vec{V}^{\infty}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)=0.$
So, for almost all $(t,s,\omega)\in\mathbb{R}^{2}\times\mathbb{S}^{2}$, we have $\mathrm{div}_{\omega}\Big{(}\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu^{t}(s,\omega)\Big{)}=0$. ∎
4.3 The Wigner measure
Proof of Theorem 3.8.
Assume Assumption 1.2 and 1.3. Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ be a solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. According to Theorem 3.7, we can consider ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$. Then, according to proof of Section 4.2.1 and Section 4.2.2, we conclude. ∎
Proof of Theorem 3.10.
Assume Assumption 1.2 and 1.3. Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ a solution to equation (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
On the one hand, according to Lemma 2.3, we know the evolution of any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ outside the curve $\mathcal{C}$.
On the other hand, according to Lemma 3.2, we consider the family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ a family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then, by applying the Lemma 4.1, we have the link between any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ over the curve $\mathcal{C}$ and the associated measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}$. Then, by applying Theorem 3.8, we have the evolution of the semiclassical measure over the curve $\mathcal{C}$.
Finally, by applying Remark 3.5 and by calculating the shape of the two-scale semiclassical measure at finite distance according to expressions (5) and (28), we conclude. ∎
4.4 Proof of Theorem 1.4 and Theorem 1.5
Proof of Theorem 1.4.
Assume Assumption 1.2 and 1.3. Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Let ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ be the solution to equation (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then, applying Theorem 3.10, according to Remark 3.5 and Corollary 3.6, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and ${\Gamma^{t}\mathrm{d}\gamma^{t}}$ a two-scale semiclassical measure at finite distance, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{tr}\left(a\big{(}\gamma(s)\big{)}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\right)\mathrm{d}t,$
with for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
We conclude thanks to the conservation of $\varepsilon$-oscillation (Lemma 2.2). ∎
Proof of Theorem 1.5.
Assume Assumption 1.2 and 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition. A consequence of Theorem 3.8 and Remark 3.5 is
$\mathrm{d}\nu\mathds{1}_{\mathcal{C}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{Tr}\Big{(}\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathcal{C}}\mathrm{d}t$
where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6). We conclude with $\mathrm{Tr}\left(\Pi_{n}(s,\sigma)\right)=1$. ∎
Appendix A Proof of compactness at infinity
Lemma A.1 (Conservation of compactness at infinity).
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1) is compact at infinity uniformly for $t\in[-T,T]$.
Proof of Lemma A.1.
Let $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ be compact at infinity. The proof relies on the analysis of
$\psi_{R}^{\varepsilon}(t)\coloneqq\chi_{R}(x)\psi_{t}^{\varepsilon},$
where ${\chi_{R}(x)=\chi\left(\frac{x}{R}\right)}$ with $\chi\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},[0,1]\right),\chi\equiv 0$ for $\left|x\right|<1$ and $\chi\equiv 1$ for $\left|x\right|>2$. We observe that $\psi_{R}^{\varepsilon}$ solves
$\mathrm{i}\varepsilon\partial_{t}\psi_{R}^{\varepsilon}=\mathrm{H}_{\varepsilon}\psi_{R}^{\varepsilon}+\varepsilon\left[\chi_{R}(x),\begin{pmatrix}0&\mathrm{D}_{1}-\mathrm{i}\mathrm{D}_{2}\\ \mathrm{D}_{1}+\mathrm{i}\mathrm{D}_{2}&0\end{pmatrix}\right]\psi_{t}^{\varepsilon}.$
Nevertheless, the first derivatives of $\chi$ are bounded therefore there exists $C>0$ such that
$\left\lVert\left[\chi_{R}(x),\begin{pmatrix}0&\mathrm{D}_{1}-\mathrm{i}\mathrm{D}_{2}\\ \mathrm{D}_{1}+\mathrm{i}\mathrm{D}_{2}&0\end{pmatrix}\right]\right\rVert_{\mathcal{L}(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right))}\leqslant\frac{C}{R}.$
Then, by an energy argument,
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\left\lVert\psi_{R}^{\varepsilon}(t)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant\frac{C}{R}\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
so
$\left\lVert\psi_{R}^{\varepsilon}(t)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $\leqslant\left\lVert\psi_{R}^{\varepsilon}(0)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\frac{C}{R}\left|t\right|\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
$\underset{R\to+\infty}{\longrightarrow}0.$
Appendix B Construction of semiclassical measures
B.1 Semiclassical measures for families valued in a separable Hilbert space
We will use the more general framework of [14] and consider families valued in a separable Hilbert space $\mathcal{H}$. More precisely, in this section, $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$. Note that in section 2.1, $\mathcal{H}=\mathbb{C}^{2}$.
Lemma B.1 (Operator valued measure).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists ${(\varepsilon_{k})_{k>0}}$ a vanishing sequence and a positive operator valued measure $\Gamma\mathrm{d}\gamma$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi),$
with $\mathcal{K}\left(\mathcal{H}\right)$ the set of compact operators on $\mathcal{H}$.
Proof of Lemma B.1.
Let $\mathcal{H}$ be a separable Hilbert space, ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, $(h_{n})_{n\in\mathbb{N}}$ a Hermitian basis of $\mathcal{H}$ and $\left(\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}\right)_{n\in\mathbb{N}}$ families of $\mathrm{L}^{2}\left(\mathbb{R}\right)$ such that $x$ a.e.
$f^{\varepsilon}(x)=\sum_{n\in\mathbb{N}}\underbrace{f^{\varepsilon}_{n}(x)}_{\in\mathbb{C}}h_{n}.$
Then the families $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ satisfy, for some constant $C>0$,
$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (36)
For all $n\in\mathbb{N}$, $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ are bounded in $\mathrm{L}^{2}\left(\mathbb{R}\right)$. By a diagonal extraction process, we can find a sequence ${(\varepsilon_{k})_{k>0}}$ and a family of Radon measures $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that for all $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}\right)$ and for all $(n,m)\in\mathbb{N}^{2}$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(b)f^{\varepsilon_{k}}_{n},f^{\varepsilon_{k}}_{m}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{tr}\Big{(}b(x,\xi)\mathrm{d}\mu_{n,m}(x,\xi)\Big{)},$
with ${\mathrm{d}\mu_{n,n}\left(\mathbb{R}^{2}\right)\leqslant\limsup_{k\to+\infty}\left\lVert f^{\varepsilon_{k}}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}}$ and $\mathrm{d}\mu_{n,m}$ absolutely continuous with respect to $\mathrm{d}\mu_{n,n}$ and $\mathrm{d}\mu_{m,m}$. Moreover, for all $(n,m)\in\mathbb{N}^{2}$, $\mathrm{d}\mu_{n,n}$ is a positive Radon measure and $\mathrm{d}\mu_{m,n}=\overline{\mathrm{d}\mu_{n,m}}$.
First of all, we construct the operator valued measure $\Gamma\mathrm{d}\gamma$. By (36),
$\limsup_{k\to+\infty}\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon_{k}}_{n}\right\rVert_{\mathrm{L}^{2}}^{2}<+\infty,$
then ${\mathrm{d}\gamma\coloneqq\sum_{n\in\mathbb{N}}\mathrm{d}\mu_{n,n}}$ defines a positive Radon measure of finite mass. By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\gamma$-measurable functions $\left(\Gamma_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
$\mathrm{d}\mu_{n,m}=\Gamma_{n,m}\mathrm{d}\gamma$
with $0\leqslant\Gamma_{n,n}\leqslant 1$ and $\Gamma_{n,m}=\overline{\Gamma_{m,n}}$ $\mathrm{d}\gamma$-almost everywhere. Let us define the operator valued function $\Gamma:(x,\xi)\mapsto\Gamma(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
$\left\langle\Gamma(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=\Gamma_{n,m}(x,\xi).$
Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $\Gamma(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}\Gamma_{n,n}(x,\xi)\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\gamma_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
Therefore, the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ generates an operator measure $\Gamma\mathrm{d}\gamma$. We now have to prove that $\Gamma\mathrm{d}\gamma$ plays the expected role.
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, $\Pi_{\leqslant n}$ the projector on the finite dimensional subspace of $\mathcal{H}$ generated by $(h_{m})_{m\leqslant n}$ and consider the finite rank operator-valued function $a_{n}\coloneqq\Pi_{\leqslant n}a\Pi_{\leqslant n}$ which belongs to $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, then
$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (37)
By passing to the limit, thanks to the construction of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}\left(a_{n}\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle$ $=\sum_{m,\ell\leqslant n}\left\langle\mathrm{Op}_{\varepsilon_{k}}\Big{(}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Big{)}f^{\varepsilon_{k}}_{m},f^{\varepsilon_{k}}_{\ell}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)}$
$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\gamma_{m,\ell}(x,\xi)$
$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Gamma_{m,\ell}\mathrm{d}\gamma(x,\xi)$
$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a_{n}(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi).$
Then, with (37) and Lebesgue dominated convergence Theorem, we have
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\big{(}a(x,\xi)-a_{n}(x,\xi)\big{)}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
B.2 Two-scale semiclassical measures
In this section, we prove Theorem 3.3.
Proof of Theorem 3.3.
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. For simplicity, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu$ of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$. We use the decomposition (17). First we analysis of the concentration of $a^{\delta}$.
  • $\bullet$
    Construction of $\mathrm{d}\mu$ outside $\mathcal{C}$
For $a\in\mathfrak{A}$ such that $a=0$ in a neighborhood of $\mathcal{C}$, then for $\varepsilon$ small enough, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y;\sigma,\eta;\frac{y}{\sqrt{\varepsilon}},\frac{\sigma}{\sqrt{\varepsilon}},\frac{\eta}{\sqrt{\varepsilon}}\right)\right)$
$=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\right).$
After extraction of a subsequence that we still denote by $(\varepsilon)_{\varepsilon>0}$, we have
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a^{\delta}\right)=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right).$
Then we construct the pair $\Gamma\mathrm{d}\gamma$.
  • $\bullet$
    Construction of $\Gamma\mathrm{d}\gamma$
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$, since $a$ is compactly supported, we recall (21)
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}+\mathcal{O}\left(\sqrt{\varepsilon}\right),$
The function $b$ is valued in the set of compact operators on $\mathcal{H}$. We can see ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ as a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We are thus doing classical calculus on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We apply Lemma B.1, then there exists $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ a vanishing sequence such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$,
$\lim_{k\to+\infty}\left\langle\mathfrak{Op}_{\varepsilon_{k}}\left(a\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0;0,0;y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Then, we apply this property to $a_{\varepsilon,R}$. Finally we construct the pair $M_{\infty}\mathrm{d}\nu_{\infty}$.
  • $\bullet$
    Construction of $M_{\infty}\mathrm{d}\nu_{\infty}$
Let $a\in\mathfrak{A}$, then we recall (20)
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Bigg{(}a(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\\ &\times\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\Bigg{)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ =&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\underbrace{\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\right)}_{\eqqcolon\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ &+\mathcal{O}\left(\delta\right).\end{split}$
therefore we consider the family $\widetilde{f}^{\varepsilon}\coloneqq\mathfrak{T}_{\sqrt{\varepsilon}}f^{\varepsilon}$ which is bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then the quantity
$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)\coloneqq\left\langle\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
is uniformly bounded in $\varepsilon>0$ and $R>1$. $a^{R}_{\delta}$ is supported at distance at least $R\sqrt{\varepsilon}$ of $\mathcal{C}$, so
$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)=\left\langle\mathrm{Op}_{1}\left(a_{\infty}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},$
as soon as $R$ is large enough, depending of $\varepsilon$. We then deduce by a diagonal extraction argument that there exists a subsequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$, a sequence $\left(R_{n}\right)_{n\in\mathbb{N}}$ and a sequence $\left(\delta_{n}\right)_{n\in\mathbb{N}}$ such that for all $a\in\mathfrak{A}$,
$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\underset{n\to+\infty}{\longrightarrow}\mathscr{I}\left(a_{\infty}\right).$
It remains to prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a measure, which will allow us to define $M_{\infty}\mathrm{d}\nu_{\infty}$. First, we prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued distribution. We observe that there exists $C>0$ such that for all $a\in\mathfrak{A}$, for all $n\in\mathbb{N}$,
$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\leqslant C\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right),$
with $\mathrm{N}_{d}$ a Schwarz semi-norm. We have $\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\to\mathrm{N}_{d}\left(a_{\infty}\right)$ then $\mathscr{I}\left(a_{\infty}\right)\leqslant C\mathrm{N}_{d}\left(a_{\infty}\right)$. Therefore $\mathscr{I}$ is a matrix-valued distribution. Secondly, we observe that the operators $a\mapsto\mathrm{Op}_{1}\left(a^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$ satisfy a semiclassical calculus in the parameters $\varepsilon$, $1/R$ and $\delta$. We have for all $a,a_{1},a_{2}\in\mathfrak{A}$,
  • $\bullet$
    Adjoint : $\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)^{\dagger}=\mathrm{Op}_{1}\left(\overline{\widetilde{a}^{R}_{\delta}}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$
  • $\bullet$
    Symbolic Calculus : as operator of $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$,
    $\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{1,\delta}\Big{)}\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{2,\delta}\Big{)}=\mathrm{Op}_{1}\Big{(}\left(\widetilde{a}_{1}\widetilde{a}_{2}\right)^{R}_{1,\delta}\Big{)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
  • $\bullet$
    Weak Gårding inequality : if $a\geqslant 0$, then for all $\widetilde{\delta}$, there exists $C_{\widetilde{\delta}}>0$, such that for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
    $\left\langle\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{\delta}\Big{)}\widetilde{f},\widetilde{f}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\widetilde{\delta}+C_{\widetilde{\delta}}\left(\varepsilon+\frac{1}{R}\right)\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
Therefore $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued measure defined on ×$\mathbb{R}\times\mathbb{S}^{2}$ that we denote $M\mathrm{d}\nu$. ∎
Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
In this appendix, we study the spectrum of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ introduced in (34).
Theorem C.1 (Spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$).
For all $(s,\sigma)\in\mathbb{R}^{2}$, the spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ consists in multiplicity one eigenvalues
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
with for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;,\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix},$
$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hilbert functions defined in (27). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ defined by
$\Pi_{0}(s,\sigma)\coloneqq g_{0}^{s}\otimes g_{0}^{s},\quad\quad\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$
are in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$.
Notably,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
Proof of Theorem C.1.
First, we will rescale the function in order to have the form $f\left(s,r(s)y\right)$. We can do it because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2). Then we dilate with $r(s)^{-1}$ and we denote by $\zeta$ the ratio $\sigma/r(s)$. For all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Bigg{[}f\Big{(}r(s)y\Big{)}\Bigg{]}$ $=\left[\begin{pmatrix}\sigma&r(s)(y+\mathrm{i}\mathrm{D}_{y})\\ r(s)(y-\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}$
$=r(s)\left[\begin{pmatrix}\zeta&y+\mathrm{i}\mathrm{D}_{y}\\ y-\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}.$
Let us define
$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}\\ \mathfrak{a}^{\dagger}&-\zeta\end{pmatrix},$
where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$. Then the link between the spectrum is
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=r(s)\mathrm{Sp}(\mathscr{T})$
So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$. Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
$\left\{\begin{matrix}2\zeta f&+&\mathfrak{a}g&=&\left(\lambda_{\zeta}+\zeta\right)f\\ &&\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}+\zeta\right)g\end{matrix}\right.$
If $\lambda_{\zeta}=-\zeta$, the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $f=0$ whence $\mathfrak{a}g=0$ so $g=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$. So $-\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}0\\ \mathfrak{h}_{0}\end{pmatrix}}$ a normalized eigenfunction. If $\lambda_{\zeta}\neq-\zeta$, then
$\left\{\begin{matrix}\mathfrak{a}\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)f\\ g&=&\frac{1}{\lambda_{\zeta}+\zeta}\mathfrak{a}^{\dagger}f\end{matrix}\right.$
and eigenvalues of $\mathfrak{a}\mathfrak{a}^{\dagger}$ are known so we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\}=\mathrm{Sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right),$
and for some $B\in\mathbb{C}$,
$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\mathfrak{h}_{|n|-1}(y)\\ \frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}+\zeta}\mathfrak{h}_{|n|}(y)\end{pmatrix}.$
Then
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes. ∎